### v1

# MIRAGE: A Benchmark for Multimodal Information-Seeking and Reasoning in Agricultural Expert-Guided Conversations

Abstract
We introduce MIRAGE, a new benchmark for multimodal expert-level reasoning and decision-making in consultative interaction settings. Designed for the agriculture domain, MIRAGE captures the full complexity of expert consultations by combining natural user queries, expert-authored responses, and image-based context, offering a high-fidelity benchmark for evaluating models on grounded reasoning, clarification strategies, and long-form generation in a real-world, knowledge-intensive domain. Grounded in over 35,000 real user-expert interactions and curated through a carefully designed multi-step pipeline, MIRAGE spans diverse crop health, pest diagnosis, and crop management scenarios. The benchmark includes more than 7,000 unique biological entities, covering plant species, pests, and diseases, making it one of the most taxonomically diverse benchmarks available for vision-language models, grounded in the real world. Unlike existing benchmarks that rely on well-specified user inputs and closed-set taxonomies, MIRAGE features underspecified, context-rich scenarios with open-world settings, requiring models to infer latent knowledge gaps, handle rare entities, and either proactively guide the interaction or respond. We evaluate more than 20 closed and open-source frontier vision-language models (VLMs), using an ensemble of reasoning language models as evaluators, highlighting the significant challenges posed by MIRAGE. Despite strong performance on conventional benchmarks, state-of-the-art VLMs struggle on MIRAGE, particularly in scenarios encountering rare entities and addressing open-ended user requests. To support model development, we fine-tune Qwen2.5-VL models on MIRAGE, observing measurable performance gains and demonstrating MIRAGE’s potential as both a benchmark and a development suite for in-domain visual reasoning and conversational decision-making in real-world settings. Our dataset 111Dataset: https://huggingface.co/MIRAGE-Benchmark and code 222Code: https://github.com/MIRAGE-Benchmark/MIRAGE-Benchmark are all publicly available. Project Page: https://mirage-benchmark.github.io/
1 Introduction
Advances in large vision–language models (LVLMs) have significantly improved AI’s ability to interpret images and generate natural-language responses. However, existing benchmarks predominantly focus on short-form visual question answering [68, 30, 12, 51], captioning [53, 50], or grounded generation under constrained contexts [29]. These tasks fall short of capturing the interactive, decision-centric nature of real-world expert consultations, where users often present open-ended, ambiguous, and visually grounded queries. In knowledge-intensive domains such as agriculture, medicine, and engineering, expert consultations inherently span multiple modalities [49] and success hinges not just on perception or language fluency, but on the ability to reason causally, handle missing context, and make interaction-level decisions. This complex interplay between multimodal understanding and contextual reasoning in professional settings represents a significant gap in the current LVLM evaluation frameworks. LLMs and LVLMs are increasingly used in domains like healthcare, law, and plant care [57, 59, 38, 71, 17, 48, 63], yet current benchmarks underrepresent knowledge-intensive scenarios involving ambiguous, multimodal queries [30]. Agriculture illustrates this gap: farmers often seek image-based, expert-level guidance during critical events, but inaccurate or unsupported model outputs [69, 40, 67] can lead to serious consequences. This highlights the need for rigorous evaluation frameworks that assess VLMs in more realistic settings.
To address this need, we introduce MIRAGE, a comprehensive benchmark designed around four core principles. First, underspecification: unlike traditional VQA and multimodal understanding datasets, MIRAGE presents user turns with latent knowledge gaps, requiring inference of missing context. Second, multimodality: each task combines natural language, images, and real-world metadata like location and time, reflecting the inputs experts receive. Third, decision-making: MIRAGE evaluates not only factual accuracy but also a model’s ability to simulate expert conversational behavior by deciding whether to ask clarifying questions or provide actionable answers. Lastly, domain grounding: built from real expert-user conversations, MIRAGE ensures ecological validity and high relevance for agricultural consultation tasks. It features problems sourced from 37,512 carefully selected high-quality user-expert conversations distilled from a corpus of 218,000 interactions collected between 2012 and 2025. Spanning more than 7000 unique biological entities (see Table 4) across plants, pests, and diseases, we evaluate models in two unique challenges absent in current benchmarks (Figure 4) A.) MIRAGE-MMST (Multimodal Single-Turn): Given a user query and associated images(s), can a model identify key biological entities, reason about causal symptoms, and produce actionable management recommendations? B.) MIRAGE-MMMT (Multimodal Multi-Turn): In an ongoing conversation, can a model decide whether to seek clarification or respond, and generate the appropriate follow-up utterance?
We evaluate 22 SOTA proprietary and open-source LVLMs covering 6 model families on MIRAGE. Our key contributions and findings are summarized below:
MIRAGE-MMST is highly challenging: Even GPT-4.1 achieves only 43.9% Identification Accuracy.
There is a pronounced disparity in performance between open-source LVLMs and GPT-4.1. The highest performing open-source models, such as Qwen2.5-VL-72B achieve approximately 29.8% in Identification accuracy and 2.47 (out of 4) reasoning score.
We introduce a novel evaluation framework using an ensemble of reasoning-capable LLMs as judges, enabling interpretable, reproducible scoring across fine-grained criteria, including Accuracy, Relevance, Completeness, and Diagnostic Parsimony.
MIRAGE exposes a substantial generalization gap: Even after LoRA fine-tuning, models like Qwen2.5-VL-3B achieve up to 28.4% accuracy on seen entities, but only 14.6% on unseen entities, revealing a persistent 14-point gap that underscores the difficulty of open-world generalization in long-tail settings.
Decision-making under partial observability of user goals remains difficult: On MIRAGE-MMMT, even top models achieve only 63% decision accuracy, with frequent errors in determining when clarification is necessary.
We believe MIRAGE will serve as a valuable benchmark for building next-generation multimodal assistants that are not only perceptually grounded but also capable of contextual reasoning, eliciting missing contextual information, and cautious recommendations in complex real-world settings.
2 The MIRAGE Benchmark
We introduce MIRAGE (Multimodal Information-seeking and Reasoning in AGricultural Expert-Guided conversations), a benchmark purpose-built to evaluate vision-language models on expert-level reasoning and decision-making in real-world in-domain consultations. It is a multimodal benchmark derived from over 200,000 raw interactions between real users and certified agronomy experts from AskExtension 333Ask Extension is an online service that connects the public with research-based answers from Cooperative Extension and university experts at U.S. Land-Grant institutions, offering guidance on agriculture, gardening, food safety, and more through a simple online form. https://ask.extension.org/. In addition to supporting studies on agricultural and knowledge-intensive LVLMs, MIRAGE uniquely assesses models across multimodal perception, causal reasoning, and clarification strategy in realistic, user-initiated scenarios.
2.1 Overview of MIRAGE
MIRAGE comprises two components: MMST, with over 29,000 single-turn expert-user interactions, and MMMT, a multi-turn corpus of 6,306 dialogues ( 12,000 turns) requiring clarify-or-respond decisions. Tasks span identification, diagnostic reasoning, and issue management guidance across seven agronomic categories (Table 4). MMST includes Standard and Contextual subsets (Section 3, Appendix C), and a fine-grained evaluation framework assesses both accuracy and utility. With over 7,600 biological entities, including thousands of unseen species in the evaluation set, MIRAGE introduces a challenging open-world generalization setting for LVLMs. Refer Appendix C for more details.
| Dataset | Type | Multimodal | Training Set | Expert Domain | Factuality | Expert Authored | Multi Turn |
| iNat21 [58] | CLS | ✗ | ✓ | ✓ | – | – | – |
| TreeOfLife [54] | CLS | ✓ | ✓ | ✓ | – | – | – |
| SimpleQA [61] | OEQ | ✗ | ✗ | ✗ | ✓ | – | – |
| MMMU [68] | MCQ | ✓ | ✗ | ✓ | ✗ | ✗ | ✗ |
| AgMMU [30] | MCQ+OEQ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ |
| CROP [70] | LFQ/Conv | ✗ | ✓ | ✓ | ✓ | ✓ | ✓ |
| MIRAGE | LFQ/Conv | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
2.2 Benchmark Comparison and Positioning
MIRAGE builds on a growing body of benchmarks in agriculture and multimodal reasoning, yet fills a critical gap left by prior efforts. Datasets like MMMU [68] focus on generalist multimodal reasoning but rely on constrained multiple-choice formats. AgMMU [30] targets agriculture, yet uses synthetic short-form MCQs that are not representative of real expert-user interactions. Domain-specific datasets such as TreeOfLife [54] and iNat21 [58] emphasize fine-grained image classification but operate in closed-world, non-interactive settings. The CROP benchmark [70] introduces multi-turn crop science QA, but is limited to two crops, lacks visual input, and is fully text-based. In contrast, MIRAGE is the only benchmark constructed from large-scale, real-world agricultural consultations and supports both multimodal single-turn and multi-turn tasks grounded in naturally underspecified user queries. As summarized in Table 1, MIRAGE includes contextual question answering, real user-submitted images with varied quality and lighting, and domain-expert-authored responses as ground truth. It spans diverse query types—including identification, causal reasoning, and management—and introduces goal-state modeling, clarify-or-respond decision policies, and a multidimensional evaluation framework that moves beyond correctness to assess identification accuracy, causal justification, response quality, and diagnostic parsimony. These features make MIRAGE a uniquely realistic and challenging testbed for evaluating LVLMs in real-world consultation scenarios.
3 MIRAGE-MMST: Multimodal Singleturn Benchmark
MIRAGE-MMST is a benchmark designed to assess expert-level, single-turn reasoning in multimodal agricultural consultations. The task setup is similar to a Long-form VQA task [37]. Each instance consists of a natural language question paired with one or more user-provided images and associated metadata (e.g., timestamp, location). Each instance consists of a natural language question , an associated image set , and metadata . Formally, a single-turn instance is represented as a triplet , and the model must generate a structured response , where denotes identified entities (e.g., crop, pest, disease), is a causal explanation, and is a management recommendation, if requested. The task evaluates the model’s ability to reason causally about visual symptoms, identify relevant agronomic entities, and, when prompted, generate detailed management recommendations grounded in the observed evidence.
To support varying levels of difficulty and contextual grounding, MIRAGE-MMST is divided into two subsets: a Standard subset, consisting of self-contained questions that can be answered using only the provided text and image, and a Contextual subset, where successful interpretation depends on implicit information such as time, location, or agricultural context not present in the input. The standard subset evaluates a model’s ability to identify entities, infer causal relationships, and generate explanatory or recommendation-based responses. In contrast, the contextual subset contains queries with latent information gaps and elliptical language, requiring models to reconstruct missing context using external priors. We first manually annotated a seed set of contextual examples and then adopted an automated classifier to separate the full dataset. See Appendix C.3 for details.
MIRAGE-MMST serves as a diagnostic tool to benchmark the capabilities of language and vision-language models in grounded agricultural reasoning. Evaluation focuses on the model’s ability to perform fine-grained entity recognition, generate causal explanations, and provide accurate and contextually appropriate management advice, when applicable.
3.1 Dataset Curation
Our dataset was collected from Ask Extension444https://ask.extension.org/, an online platform where users submit agricultural queries and receive expert responses from trained volunteers and professionals. We aggregated 218K single-turn dialogues from December 2012 to April 2025. To ensure high dataset quality and relevance, we implemented a rigorous four-step curation process as shown in Figure 5 (1) Data Cleaning, removing incomplete or unsuitable dialogues; (2) Data Categorization, organizing dialogues into two primary subsets, Subset 1: Identification (Plant Identification, Insect and Pest Identification, Plant Disease Identification) and Subset 2: Management (Plant Disease Management, Insect and Pest Management, Plant Care and Gardening Guidance, Weeds/Invasive Plants Management), extracting and collecting main entity name, and labeling context-aware responses; (3) Data Reformatting, involving the removal of personal information, enhancing questions with supplementary content from expert-provided URLs using GPT-4.1, and reconstructing Identification task answers with visual enhancements through GPT-4.1-mini to ensure consistent formatting, detailed visual descriptions, and comprehensive reasoning chains; and (4) Data Splitting, partitioning the curated dataset into Standard Benchmark (8,184 dialogues), Standard Training (17,532 dialogues), and Contextual Benchmark (3,934 dialogues) based on contextual labeling. Further curation details are provided in the Appendix C.2.
4 MIRAGE-MMMT: Multimodal Multiturn Benchmark
MIRAGE-MMMT is a multimodal decision-making task, grounded in real-world agricultural consultations. Users pose complex, often image-supported questions about plant health, pest identification, growing conditions, and other agronomic concerns. Each dialogue reflects a practical scenario in which the expert must reason over conversation history and visual context to decide: (1) whether to respond with guidance based on what is known, or (2) whether to pause and seek additional input to resolve a knowledge gap. This introduces a decision-making challenge tightly coupled with natural language generation.
Task Formulation:
We formalize this as a joint decision making-generation task: Given a multi-turn dialogue , where and are utterances, and an associated image set , the model must jointly predict a decision and generate the corresponding utterance .
The model infers the user’s goal and a goal state . It selects:
Then it generates:
The model is evaluated on decision quality and generation utility via LLM-as-a-Judge.
Domain-General Task Modeling Framework: While MIRAGE-MMMT is instantiated in the agricultural domain, the underlying task formulation, where an expert must decide between seeking clarification or issuing a response based on evolving multimodal dialogue context, generalizes to a broad class of consultative, decision-oriented interactions. This includes domains such as healthcare triage, customer support, legal advising, technical troubleshooting, and educational tutoring. Crucially, our framework is dataset-agnostic: given any multi-turn consultation corpus with follow-up user inputs (text and/or images), our pipeline can be applied to generate structured <Clarify> or <Respond> training data.
4.1 Dataset Curation
The goal of MIRAGE-MMMT is to model the decision-making process of an expert who must decide, based on the evolving multimodal context, whether to respond with actionable guidance or clarify by seeking more information. To simulate this, we extract segments of each dialogue such that the input context ends with a user utterance, and a later user message, typically written in response to an earlier expert query, is treated as a revealed fact. This allows us to reconstruct the decision point at which the expert, based on available context and images, would determine whether the current information is sufficient to respond or whether a clarification is necessary. We prompt a powerful LLM (e.g., GPT-4o) to generate a structured task instance consisting of dialogue context, referenced images, and metadata such as geographic location and topic. To ensure data safety, we perform automated PII sanitization while preserving domain relevance. (see Appendix F.2 for more details)
5 Experiments
We evaluate a diverse set of models on both MIRAGE-MMST and MIRAGE-MMMT benchmarks to assess expert-level multimodal reasoning as well as their ability to make accurate clarify-or-respond decisions and generate goal-consistent, grounded responses.
5.1 Model Selection
The LVLMs we evaluated can be divided into two groups: 1.) Proprietary Models: this group comprises SOTA GPT models (4.1, 4.1-mini, 4o, 4o-mini) [7], Claude-3 (3.5v2 Sonnet, 3.7 Sonnet, 3 Haiku) [11] available through API service. 2.) Open-Source Models: this group includes LLaMa-4 Scout-17B [44], LLaVa (7B mistral, 7B qwen2) [41], Qwen 2.5 VL models (3B, 7B, 32B, 72B) [5], Gemma 3 models (4b, 12, 27b) [56] and Intern VL3 (2B, 8B, 14B, 38B, 72B) [73]. 3.) Finetuned Models We perform multi-image multimodal fine-tuning of the Qwen 2.5 VL models on the MMST dataset. All experiments were conducted with NVIDIA A100-40GB/H200-141GB GPUs.
5.2 Evaluation Methodology
To facilitate consistent, interpretable, and reproducible evaluation of model outputs, we implement a structured LLM-as-a-Judge protocol. Rather than relying on single-model, single-pass assessments, we leverage an ensemble of reasoning-capable language models: DeepSeek-R1-Distilled [34], Qwen3-32B [5], and Phi-4-Reasoning [6], as judges. Each candidate’s response is evaluated across three independent generations, with every generation scored by the full ensemble, resulting in nine total evaluations per sample. This design captures both linguistic variability across generations and scoring consistency across evaluators. To validate the fidelity of this framework, we quantify inter-rater agreement using Fleiss’ for categorical scoring and Kendall’s W for ordinal coherence, ensuring both statistical reliability and scoring consistency across the ensemble. Fleiss’ scores for the ID task (binary classification) generally fall within the 0.75–0.88 range, indicating “good” to “excellent” agreement by standard interpretation guidelines. Full details, scoring rubrics, and agreement statistics are provided in Appendix D.1.1. The prompt for evaluation is showed in Figure 14.
Figure 30 illustrates our evaluation framework in action. The judge model is given the expert-authored gold response and systematically evaluates both the correctness of the model’s prediction and the quality of its explanation. By reasoning step-by-step, the judge produces an interpretable “judgment trace” that makes the evaluation process transparent and highlights where the model’s output aligns with, or diverges from, the expert’s reasoning.
5.3 MIRAGE-MMST Results
We evaluate model performance on the MMST-ID task using two complementary metrics: 1.) Identification Accuracy which is a binary metric that measures whether the entity identified by the model matches the expert’s answer. A response is scored as correct (1) if the predicted entity string exactly matches any of the reference fields: entity name, scientific name, or common names; otherwise, it receives a score of 0. Reasoning Score, evaluates the quality of the model’s visual and linguistic justification for its prediction. It is graded on a 0–4 scale by the judges, based on the presence of key visual clues, descriptive specificity, and causal coherence. Higher scores reflect more complete, interpretable, and expert-aligned justifications.
To evaluate model outputs on the MMST-MG task, we adopt a multidimensional scoring framework that assesses both factual correctness and communication quality. Each model response is rated on a 0–4 scale across the following four dimensions the judges: 1.) Accuracy: which measures the factual alignment with the expert’s response. A perfect score (4) is awarded when the facts, terminologies, and recommendations fully align with the expert response. Relevance measures the ability of the model to stay on-topic with the user’s original query and avoid tangential information. Completeness assesses whether the model covers all key information provided by the expert, and Parsimony, inspired by Occam’s Razor; this score captures the conciseness, clarity, and actionability of the model response. We aggregate these scores in a 2:1:1:1 ratio. See Appendix D.1.2 for more details.
5.4 Performance Comparison
MMST-ID: Table2 presents the results of zero-shot prompting across 22 LVLMs on the MMST-ID task. All models are evaluated without task-specific fine-tuning, allowing for a fair comparison of out-of-the-box reasoning and grounding capabilities. We observe a consistent gap between closed-source and open-source models. GPT-4.1 achieves the highest Average Identification Accuracy (43.9%) and the top Reasoning Score (3.0), with other proprietary models like Claude 3 and GPT-4o performing competitively. In contrast, the best open-source model, Qwen2.5-VL-72B, reaches only 29.83% accuracy and 2.47 reasoning score.
MMST-MG: For cross-model comparisons, we compute a weighted aggregate score over the four evaluation metrics. Table 2 and Figure 2 indicate GPT-4.1 consistently outperforms all models across judges and metrics, achieving the highest scores in Accuracy (3.24) and Parsimony (3.01). Among open-source models, Qwen2.5-VL-32B and Gemma-3-27B perform best.
| Model | MMST-ID Task | MMST-MG Task | |||||
| Acc (%) | Reason | Acc | Rel | Comp | Pars | W-Sum | |
| gpt-4.1 | 43.9 | 3.01 | 3.24 | 3.60 | 3.22 | 3.01 | 0.82 |
| gpt-4.1-mini | |||||||
| gpt-4o | |||||||
| gpt-4o-mini | |||||||
| claude-3-7-sonnet | |||||||
| claude-3-5-sonnet | |||||||
| claude-3-haiku | |||||||
| Llama-4-Scout-17B-16E-Instruct | |||||||
| llava-v1.6-mistral-7b-hf | |||||||
| llava-onevision-qwen2-7b-ov-hf | |||||||
| Qwen2.5-VL-3B-Instruct | |||||||
| Qwen2.5-VL-7B-Instruct | |||||||
| Qwen2.5-VL-32B-Instruct | |||||||
| Qwen2.5-VL-72B-Instruct | |||||||
| gemma-3-4b-it | |||||||
| gemma-3-12b-it | |||||||
| gemma-3-27b-it | |||||||
| InternVL3-2B | |||||||
| InternVL3-8B | |||||||
| InternVL3-14B | |||||||
| InternVL3-38B | |||||||
| InternVL3-78B | |||||||
Scores are averaged over the three judge models. Closed-source rows are red/light-red, open-source rows blue/light-blue. Bold purple numbers mark the overall best for each metric.
Performance improves steadily with model scale within families, especially for open models as seen in Figure 9. However, scaling gains are not uniform across families and show early saturation in some cases. Qwen2.5-VL series show the strongest scaling behavior, Gemma and InternVL3 also improve with size, but lag behind Qwen, particularly in identification accuracy, suggesting weaker visual grounding. None of the open models approach GPT-4.1 performance, which retains a sizable lead in both accuracy and reasoning. Notably, reasoning improvements appear to saturate earlier ( 32B), indicating that scale alone may not bridge the gap without better visual-linguistic alignment or task supervision. Open-source models still lag by 0.1–0.15 in weighted score compared to GPT-4.1, especially due to verbosity (low parsimony) and subtle factual gaps (accuracy).
Model Finetuning: We utilize the MMST training dataset to fine-tune Qwen 2.5 VL models (3B, 7B, 32B) and evaluate performance on both MMST-ID & MMST-MG tasks. Figure 1 illustrates the effects of LoRA-based fine-tuning on the Qwen2.5-VL-3B model across five checkpoints. Our results indicate consistent improvements in seen entity accuracy on MMST-ID after fine-tuning, rising from 22.3% (base-Instruct) to a peak of 28.4% (epoch 6). Across model sizes, as seen in Figure 10, we see a consistent convergence behavior across epochs, with 32B achieving the strongest peak performance (37.6% at epoch 6) and reasoning score (3.04). However, models struggle to generalize, as accuracy on unseen entities remains stagnant. On the MMST-MG task, from the base-Instruct model to LoRA-ep6, we observe steady improvement across all four metrics as well. Relevance is the highest-scoring metric throughout, indicating the model’s ability to remain on-topic is strong even in zero-shot, and improves further with tuning. Parsimony also improves consistently, indicating models can learn concise delivery of recommendations.
Performance on Contextual Subset: The contextual subset of MIRAGE-MMST reveals that inferring latent context remains a major bottleneck for current LVLMs. Even top-tier models like GPT-4.1 perform well but not well, and the gap to open-source models is significant. Compared to the standard subset, scores are consistently lower across the board, indicating that reasoning under partial observability and implicit context reconstruction are significant failure points for current models, as seen in Table 8. We also observe that current LVLMs make limited use of metadata, such as location and time, despite their importance in real-world agricultural reasoning see Table 10 for comparison with and without the meta-data. Across both proprietary and open-source models, the inclusion of metadata yields minimal improvements in identification accuracy (e.g., +1.6% for GPT-4.1, +0.6% for Qwen2.5-VL-72B) and virtually no gains in reasoning scores. In some cases, performance even degrades slightly, suggesting that models may treat metadata as irrelevant or distracting.
5.5 MIRAGE-MMMT Results
Before presenting experimental results, we first outline the structure of the decision task and its dependence on input observability. Under full observability, where the model is given access to an oracle goal state that explicitly enumerates known and missing information, a rule-based oracle achieves 98.45% decision accuracy. This illustrates that the challenge lies in the model’s ability to infer missing context from dialogue history. In real-world settings, such goal states are unavailable at inference time. MIRAGE is therefore designed to evaluate models under partial observability, where decisions must be made using only the dialogue history. To study the effect of structured supervision, we trained logistic regression classifiers with progressively enriched inputs (Table 23). Performance improves modestly with access to the user-stated goal (69.79% → 71.34%), and substantially with access to the oracle goal state (89.27%), confirming that inferring implicit context is the key difficulty in this task.
5.6 Performance Comparison
We evaluate the model’s high-level decision and its alignment with user goals in three complementary ways. First, Decision Accuracy captures how often the model’s chosen action, whether to ask a clarification question or to respond directly, matches the gold annotation; To ensure that when the model generates an utterance, it is doing so meaningfully, we measure Clarify Relevance: for every turn labeled <Clarify> or Respond Relevance for all turns labeled <Respond> using the LLM judges.
| Model | Zero-Shot | CoT | (CoT–ZS) | ||||||
| Acc% | Clarify | Respond | Acc% | Clarify | Respond | Acc | Clarify | Respond | |
| GPT-4o | 65.50 | 72.80 | 78.50 | ||||||
| Claude 3.7 Sonnet | |||||||||
| LLaMA 4 Maverick 17B | |||||||||
| LLaMA 4 Scout 17B | |||||||||
| LLaMA 3.2 90B | |||||||||
| Qwen 72B | |||||||||
We conduct experiments under Zero-Shot and COT [60] prompt settings on MMMT shown in Table 3. In the zero-shot setting, GPT-4o achieves the highest Decision Accuracy (62.98%) and leads in both Clarify Goal Relevance (70.75%) and Respond Goal Relevance (77.07%), showing strong overall reasoning and generation alignment. Open models like LLaMA 4 Maverick and LLaMA 4 Scout perform competitively, reaching 53–49% decision accuracy. Introducing chain-of-thought (CoT) reasoning consistently improves performance across models. GPT-4o gains +2.5% in Decision Accuracy (to 65.5%) and improves goal relevance in both Clarify (+2%) and Respond (+1.5%) settings. Other models, especially Claude 3.7 and Qwen 72B, benefit more substantially (improvements in clarify goal relevance by +10.5%).
6 Conclusion
We introduce MIRAGE, a high-fidelity benchmark for evaluating vision-language models (VLMs) in expert-level agricultural consultations. MIRAGE addresses key limitations of existing benchmarks by incorporating real-world, multimodal, and context-rich interactions grounded in domain-specific decision-making. Despite these strengths, MIRAGE has a few limitations. It does not yet simulate interactive dialogue with real users or user simulators, limiting evaluation of adaptation and dialogue flow over time. Visual follow-ups are also not modeled; user turns beyond the initial user turn are assumed to be text-only. Looking ahead, MIRAGE opens several promising directions for advancing multimodal reasoning and conversational capabilities of models in knowledge-intensive domains. Improving performance on unseen entities will require better open-world generalization strategies, such as retrieval-augmented generation, compositional reasoning, or domain-adaptive pretraining. Limited gains from metadata highlight the need for explicit context modeling—including structured encodings of time, location, and integration with external knowledge bases capturing seasonal and biological priors. Persistent challenges in clarify-or-respond decision-making point to a need for models that can infer latent user goals and reason about missing information. We hope this benchmark enables the development of more context-sensitive, knowledge-intensive VLMs. MIRAGE is publicly released to support the development of vision-language systems that go beyond narrow question answering, toward models capable of engaging in natural interactions that involve ambiguity, a need for clarification, visual understanding, and decision-making when critical context is implicit or missing, addressing common challenges in real-world expert consultations.
7 Limitations
While MIRAGE provides a comprehensive evaluation framework for multimodal expert-level reasoning within agricultural consultations, several inherent limitations remain. First, the scope of the benchmark primarily reflects smallholder and backyard gardening scenarios, sourced largely from AskExtension’s database. Consequently, it does not fully encompass large-scale industrial agricultural practices or the more complex logistical, economic, and environmental issues prevalent in commercial farming operations. Additionally, agriculture is an extensive and multidisciplinary domain, covering areas such as soil science, agricultural engineering, economics, environmental sustainability, and policy implications. The current dataset captures only a subset of these dimensions. In the multi-turn scenario (MMMT), although designed to assess models’ ability to reason over dialogue context, the task does not yet simulate dynamic, truly interactive exchanges between models and users. The evaluations focus primarily on the models’ capacity to interpret and infer latent contextual information and to determine appropriate conversational actions based on partial observability. Consequently, MIRAGE does not fully measure interactive conversational reasoning, such as handling conversational repairs or real-time user feedback.
8 Impact Statement
MIRAGE introduces a high-fidelity benchmark that addresses a critical gap in evaluating vision-language models (VLMs) for real-world, expert-level decision-making in knowledge-intensive domains. By grounding tasks in over 35,000 authentic agricultural consultations involving multimodal, context-rich queries, MIRAGE pushes the boundaries of current LVLM capabilities beyond perception and language generation to include causal reasoning, clarification strategies, and open-world generalization. It provides both a rigorous testbed and a development suite for building safer, more reliable AI assistants in domains where decision quality has tangible real-world consequences. However, caution is advised in fully relying on MIRAGE results: the benchmark, while comprehensive, does not yet simulate dynamic user interactions or evolving contexts, and agriculture itself encompasses a vast range of specialized knowledge not fully captured in this dataset. Thus, models performing well here may still face significant challenges in broader, real-world agricultural deployments, particularly in novel or unforeseen scenarios.
9 Acknowledgments
This work was supported by the Amazon-Illinois Center on AI for Interactive Conversational Experiences Award, the AIFARMS National AI Institute, and the Center for Digital Agriculture at the University of Illinois. We are grateful to the AskExtension team for providing the dataset and to Dr. John F. Reid, Dennis Bowman, and Talon Becker for contributing their domain expertise and invaluable feedback. The contributions of Shubham Garg and Hooshang Nayyeri to this work are independent of their roles at Amazon. The views expressed and conclusions drawn herein are solely those of the authors and do not necessarily reflect the official policies or positions, expressed or implied, of any sponsoring organizations. This work used Delta advanced computing and data resource at University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications through allocation CIS250434 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
Supplementary Material
[…]
Appendix A Data Source
The MIRAGE benchmark is constructed from a large-scale archive of real-world agricultural consultations obtained from Ask Extension [25], a U.S. national digital platform maintained by the Extension Foundation. Ask Extension is part of the broader Cooperative Extension System, a federally supported network of land-grant universities that delivers science-based, community-oriented education and services across the United States. The platform connects members of the public, such as farmers, gardeners, or homeowners, with university-affiliated experts who provide timely, research-backed responses to their questions.
Inquiries submitted through the Ask Extension portal are answered by a diverse pool of domain specialists, including university faculty, Extension educators, and trained volunteers such as Master Gardeners. These responses reflect both academic rigor and region-specific expertise, leveraging a unique model of public scholarship that blends localized agricultural knowledge with the latest findings from land-grant institutions. This institutional provenance ensures that the answers used in our dataset are highly reliable, authored by qualified experts, and grounded in scientifically validated practices.
We collected approximately 285,393 interactions (218,431 for single-turn; 66,962 for multi-turn) from the Ask Extension platform, spanning from December 2012 to April 2025. Each entry captures a real question from a user, along with the corresponding expert response, and may include user-uploaded images, time of submission, and geographic metadata.
As seen in Figure 3, Our filtered AskExtension dialogues are strongly multimodal and reference-rich. Users typically include one to three images in their questions, about 35% of turns have a single image, 26% include two, and 39% include three. Experts in turn ground their advice in external sources: roughly 70% of answers cite a single URL, while the remaining 30% provide two. The total amount of content fetched from those links spans from about 2 000 up to 10 000 tokens per response, indicating that experts draw on substantial external context to support their guidance.
Appendix B Related Works
Multimodal Large Language Models: Recent advances in multimodal large language models (LLMs) have markedly expanded vision–language reasoning capabilities. Proprietary models such as GPT-4 [8], Claude 3 Sonnet [10], and Google’s Gemini [55] demonstrate strong capabilities in unifying visual and textual modalities, achieving notable success across diverse multimodal benchmarks. Concurrently, open-source models—including Qwen-VL 2.5 [14], Gemma 3 [56], and InternVL-3 [73]—have narrowed the performance gap while remaining publicly accessible. Although these models excel on general-domain benchmarks, they underperform in agriculture: they lack fine-grained visual expertise, agronomic terminology, and the ability to reason about rare biological entities and management practices. MIRAGE is designed to expose these weaknesses by providing domain-specific, multimodal tasks that require expert-level diagnosis and advice.
Multimodal Agricultural Benchmarks: Agricultural benchmarks [19, 26, 31] have progressed from controlled, image-only datasets to more realistic vision–language corpora. Early resources such as PlantVillage [36] and PlantDoc [52] emphasise leaf-level disease classification under laboratory and in-field conditions, respectively. The AgMMU [30] benchmark augments field images with farmer–expert dialogues, yet its evaluation is dominated by short, synthetic multiple-choice questions that limit open-ended reasoning. The recent Crop Disease Domain Multimodal (CDDM) [42] corpus contributes 137k crop–disease images paired with 1 M single-turn QA pairs, but it focuses narrowly on disease identification and management. In contrast, MIRAGE is derived from large-scale, real-world consultations covering more than 7000 biological entities; it offers both single-turn and multi-turn tasks, explicitly models clarify-or-respond decisions, and scores long-form answers along multiple dimensions of reasoning quality, thereby providing a far more ecologically valid and challenging test bed.
Long-Form Question Answering Evaluation: Lexical-overlap metrics such as ROUGE [39], BLEU [46], and BERTScore [72] are ill-suited to evaluating open-ended, knowledge-intensive answers, as they correlate poorly with human judgements. Recent work therefore adopts LLM-as-Judge paradigms, exemplified by AlpacaEval [22], MT-Bench [13], and G-Eval [43], in which powerful LLMs are prompted to assess responses along axes such as factuality, coherence, and completeness. While these methods better align with human preferences, single-model evaluations remain vulnerable to bias and opacity. MIRAGE employs an interpretable ensemble of reasoning-focused LLM judges DeepSeek-R1-Distilled [34], Qwen 3-32B [5], and Phi-4-Reasoning [6] and by conducting three independent generation–evaluation passes per sample. This multi-model, multi-run protocol enhances robustness, enables variance analysis, and provides publicly inspectable rationales, yielding transparent and reproducible long-form QA evaluation.
State Tracking in Dialogue Systems: State tracking has been a central component in task-oriented dialogue systems [16], where it plays a critical role in maintaining a representation of user intent and progress toward the user’s goal throughout the interaction. Dialogue state tracking (DST) typically involves predicting a belief state, a structured representation of slot-value pairs reflecting the current user intent at every turn. The widespread adoption of state tracking in dialogue systems has enabled robust multi-turn interactions, improved task success rates, and enhanced user satisfaction. Benchmarks like MultiWOZ [15, 24, 66] and Schema-Guided Dialogue (SGD) [47] have been instrumental in shaping the field by introducing multi-domain, open-schema challenges that pushed models to become more generalizable and scalable. These successes in dialogue suggest that structured, continuously updated representations of task progress can be highly beneficial in other interactive or sequential decision-making settings. Building on this insight, we introduce Goal State Tracking (GST), a mechanism that explicitly monitors whether all user-provided information is sufficient to generate a sound consultative recommendation; when gaps remain, GST triggers the agent to pose targeted clarifying questions, ensuring that advice is delivered only once the requisite context is complete.
Appendix C MIRAGE-MMST
C.1 Benchmark Details
We summarize the key statistics of MIRAGE-MMST dataset in Table 4, which forms the single-turn evaluation component of the MIRAGE benchmark. The dataset is partitioned into three subsets: Standard, Contextual, and Standard Training Data, each designed to support different evaluation and training objectives.
| Overall Statistics | Standard | Contextual | Standard Training Data |
| Total Samples | 8 184 | 3 934 | 17 532 |
| Total Images | 15 069 | 8 069 | 33 120 |
| Per-Sample Statistics | |||
| Avg. Question Words | 69.57 | 80.94 | 67.53 |
| Avg. Answer Words | 163.13 | 222.97 | 171.18 |
| Avg. Number of Images | 1.84 | 2.05 | 1.89 |
| Category Statistics | |||
| Total Identification Tasks | 4 324 | - | 7 398 |
| Plant Identification | 2 600 | - | 3 919 |
| Insect and Pest Identification | 1 146 | - | 2 131 |
| Plant Disease Identification | 578 | - | 1 348 |
| Total Management Tasks | 3 860 | 3 934 | 8 957 |
| Plant Care and Gardening Guidance | 1 609 | 1 797 | 3 707 |
| Insect and Pest Management | 725 | 641 | 1 689 |
| Plant Disease Management | 1 047 | 1 184 | 2 445 |
| Weeds / Invasive Plants Management | 479 | 312 | 1 116 |
| Others | - | - | 1 177 |
| Entity Statistics | |||
| Plant Entities | 4 485 | 999 | 1 725 |
| Insect / Pest Entities | 1 732 | 306 | 840 |
| Plant Disease Entities | 639 | 200 | 320 |
Overall Composition: The dataset comprises a total of 29,650 high-quality user-expert interactions and over 56,000 user-submitted images. The Standard Benchmark subset contains 8,184 samples, while the Contextual Benchmark includes 3,934 samples that explicitly rely on implicit context (e.g., location-related, timing-related) not explicitly derivable from image or user query and metadata. An additional 17,532 samples are used for training and pre-tuning models.
Per-Sample Characteristics: Each sample includes both user and expert turns along with image inputs. On average, standard samples contain 69.6 words in user questions, 163.1 words in expert answers, and 1.84 images per sample. Contextual samples tend to be longer and more detailed, with 80.9 words per question, 222.9 words per answer, and a slightly higher image count (2.05 images per sample), reflecting the increased reasoning burden in these settings.
Task Category Distribution: MIRAGE-MMST covers a broad spectrum of expert tasks, which are grouped into two high-level categories: A.) Identification Tasks (7398) including, Plant Identification (3,919), Insect and Pest Identification (2,131), Plant Disease Identification (1,348) & B.) Management Tasks (8957) spanning, Plant Care and Gardening Guidance (3,707), Insect and Pest Management (1,689), Plant Disease Management (2,445), Weed and Invasive Plant Management (1,116)
Entity Coverage: The dataset includes over 7,000 unique biological entities, with fine-grained coverage across: 4,485 plant species, 1,732 insect/pest categories and 639 plant disease types.
C.2 Data Curation Details
Our benchmark utilizes real-world dialogue data sourced from online platforms, necessitating a rigorous multi-step curation process to ensure the dataset’s high quality and relevance. The curation process comprises four main steps (See Figure 5):
Step 1: Data Cleaning
The initial step involved removing unsatisfactory or incomplete data points. We first sampled the data and conducted manual checking, identifying four primary issues. Specifically, we excluded dialogues where the expert: (1) recommended contacting another individual or organization; (2) requested additional information from the user without providing a complete standalone response, as the benchmark focuses exclusively on single-turn dialogues; (3) expressed uncertainty regarding their response; or (4) explicitly indicated an inability to assist. Subsequently, we employed GPT-4o-mini to automate the filtering of these identified issues. Manual verification and automated filtering were iteratively performed to refine and optimize the filtering prompts. Final results of this process are described in Section C.3. Additionally, expert responses containing inaccessible or broken URLs were removed to ensure the dataset’s integrity and usability.
Let the initial dataset of raw expert-user dialogues be denoted as:
where is the user query, is the expert response, is associated metadata (e.g., images, timestamp, location), and is the number of raw entries.
Filtering Unvalid Samples
We define a filtering function , which retains a dialogue only if the response:
is complete (not asking for follow-up or deferring);
does not express uncertainty or redirect the user;
is not a broken or inaccessible URL;
The cleaned dataset is:
Step 2: Data Categorization
Data Classification: We categorized the dataset into subsets relevant to agricultural language model research, resulting in seven primary categories grouped into two subsets: Subset 1: Identification (Plant Identification, Insect and Pest Identification, Plant Disease Identification) and Subset 2: Management (Plant Disease Management, Insect and Pest Management, Plant Care and Gardening Guidance, Weeds/Invasive Plants Management). Dialogues not fitting these categories were labeled as "Others." We used GPT-4.1 to classify the dataset.
Let be a classifier mapping each sample to one of predefined agronomic categories. Define:
Entity Extraction: We extracted relevant entities based on the assigned category: plant names for Plant Identification, Plant Care and Gardening Guidance, and Weeds/Invasive Plants Management; insect or pest names for Insect and Pest Identification and Management; and disease names for Plant Disease Identification and Management. We used GPT-4.1-mini to extract the entities. These entities are then enriched with their synonyms (See Section C.4).
Let be entity extractors for respective domains. Then:
Contextual-Aware Labeling: Contextual-aware labeling involved analyzing expert answers that leveraged implicit context such as user location and timing. These instances included cases where experts cited location-specific regulations or practices, provided location-dependent advice, referenced current weather conditions specific to the user’s location and timing (e.g., recent drought, frost conditions), or offered time-specific recommendations. These were labeled as "location_related" or "time_related" data. We used GPT-4o and GPT-4o-mini to label the data. This step also involved manual checking. Final results of this process are described in Section C.3.
Let indicate whether a sample includes contextual elements (e.g., location-specific recommendations):
Split the dataset as:
Step 3: Data Reformatting
Content Removal and Question Enhancement: We reformatted the remaining data to enhance clarity and appropriateness for language model benchmarks. Specifically, we removed personal identification information, references specific to the "Ask Extension" service, and content unsuitable for interactions with language models (e.g., mentions of voicemails). Additionally, relevant details from dialogue titles were merged into questions when they provided additional context. We used GPT-4.1-mini model to reformat the data.
Integration of Supplementary Content: Approximately half of the expert responses contained URLs referencing supplementary information. To enrich these responses, we crawled and extracted the content from these URLs. Subsequently, we used the GPT-4.1 model to integrate this supplementary content with the original expert answers, producing more detailed and comprehensive responses.
Identification Answer Reconstruction with Visual Enhancement: Initial identification (ID) dataset expert responses included reasoning processes and conclusions but lacked consistent formatting, comprehensive visual descriptions, and complete reasoning chains. We utilized the GPT-4.1-mini model, capitalizing on its multimodal and information integration capabilities, to reconstruct and enhance these responses. This involved adding detailed descriptions of key visual characteristics (such as distinguishing features of plants and insects, or observable symptoms of plant diseases) and constructing clear, coherent reasoning chains. The standardized answers were formatted into a concise single-paragraph structure, clearly presenting both the reasoning process and the final result, thereby facilitating efficient benchmark evaluation.
Let with Plant ID, Pest ID, Disease ID. We define an enhanced answer:
The reconstructed dataset is:
Step 4: Data Splitting
We first divided the dataset into standard and contextual subsets based on contextual-aware labeling.
Standard Data: The standard data subset was partitioned into benchmark (30%) and training (70%) datasets. The splitting aimed to maximize diversity by ensuring at least one sample per entity in the benchmark, which led to some entities appearing exclusively in the benchmark due to limited samples. Additionally, we preserved the original distribution across all seven categories and maintained the proportion of URL-based responses within each category. Ultimately, we obtained a Standard Benchmark dataset of 8,184 dialogues and a training dataset of 17,532 dialogues.
Contextual Data: Context-sensitive questions were augmented with explicit user location and timing details. Due to their complexity, these were entirely allocated to the Contextual Benchmark, totaling 3,934 dialogues.
We define a stratified sampling function for selecting 30% of data for benchmarking:
C.3 Manual Check and Model Selection for Data Sanitation
As part of our data curation pipeline, we conducted a targeted analysis of user-expert conversations to identify specific characteristics that affect data quality. This step was essential for filtering low-value interactions and retaining samples that reflect sensitivity to latent contextual reasoning typical in real-world agricultural consultations.
We focused on three key characteristics:
Unsatisfactory: These include expert replies that fail to provide meaningful or actionable guidance. Common patterns include vague disclaimers such as “I’m not sure how to help you with this” or deferrals to third-party support (“You may want to contact your local extension office”). These represent non-informative speech acts and were marked for exclusion from the curated dataset to maintain high informational integrity.
Location-related: Many expert responses assumed the user’s geographic context, such as referencing local regulations, climate patterns, or soil characteristics, without this context being explicitly provided by the user. While this introduces contextual elision, these responses are not deficiencies; they reflect realistic, situated expertise. We retained these samples, recognizing their value in evaluating models’ ability to interpret or recover latent geographic/location-related context.
Time-related: Similarly, several expert responses implicitly relied on temporal context, such as seasonal crop cycles or pest development stages. These exhibited temporal underspecification, where the meaning of the advice depends on when the consultation occurred (e.g., “The pest is likely in its larval stage right now” during a spring consultation). These interactions were preserved as they reflect authentic domain-specific reasoning under temporal constraints.
To identify these characteristics at scale, we first manually annotated a stratified sample of 111 conversations. Each was labeled with one or more of the above characteristics. We then used few-shot prompting to evaluate a set of large language models (LLMs) on their ability to classify these characteristics. For each model, we provided a set of illustrative examples demonstrating how each characteristic manifests in expert replies. We measured performance using standard classification metrics: accuracy, precision, recall, and F1 score, broken down by characteristic type. (See Table 5)
| Characteristic | Count | Model | Accuracy | Precision | Recall | F1 Score |
| Unsatisfactory | 51 | gpt-4o-mini | ||||
| gpt-4o | ||||||
| gpt-4o-mini + analysis | 0.9176 | |||||
| gpt-4o + analysis | ||||||
| gemini-2.0-flash | ||||||
| gemini-2.0-flash + analysis | ||||||
| Location-related | 25 | gpt-4o-mini | ||||
| gpt-4o | ||||||
| gpt-4o-mini + analysis | 0.8400 | |||||
| gpt-4o + analysis | ||||||
| gemini-2.0-flash | ||||||
| gemini-2.0-flash + analysis | ||||||
| Time-sensitive | 16 | gpt-4o-mini | ||||
| gpt-4o | ||||||
| gpt-4o-mini + analysis | ||||||
| gpt-4o + analysis | 0.8750 | |||||
| gemini-2.0-flash | ||||||
| gemini-2.0-flash + analysis |
Based on this evaluation, we selected gpt-4o-min plus analysis for filtering unsatisfactory data and labeling location-related data; gpt-4o plus analysis for labeling time-related data in our data curation workflow.
C.4 Biological Entity Synonymy Collection
To facilitate fair evaluation of biological entity identification in model outputs, we developed a comprehensive name collection pipeline that aggregates all valid references, both scientific and vernacular, for biological entities in our dataset. Rather than standardizing to a single canonical form, our objective was to compile an exhaustive synonymy for each entity, enabling robust matching regardless of the nomenclatural variant used. The system queries multiple authoritative taxonomic databases via their APIs in a hierarchical approach, beginning with the Global Biodiversity Information Facility (GBIF) [2] and extending to iNaturalist [58] Encyclopedia of Life (EOL) [1], Integrated Taxonomic Information System (ITIS) [3], Wikipedia [62], and NCBI Taxonomy [4]. Queries are enhanced with category-specific context (e.g., kingdom Plantae for botanical entries) to improve retrieval accuracy. For each entity, we preserve all retrieved scientific names (including accepted names, synonyms, and historical nomenclature) and vernacular names across languages, with metadata indicating the source authority. This comprehensive approach prevents unfair penalization during model evaluation when, for instance, a model correctly identifies an organism using its scientific name (Phytolacca americana) while the reference answer uses a common name ("pokeweed"), or vice versa. The resulting enriched dataset maintains the original hierarchical structure while appending all valid nomenclatural alternatives, thereby supporting more equitable assessment of biological entity recognition capabilities.
Appendix D MMST Evaluation Criteria
D.1 Reasoning LLM-as-Judge
D.1.1 Traditional LLM-as-Judge
Traditional metrics for evaluating long-form question answering have significant limitations. Research by Xu et al. [64] reveals that established approaches like ROUGE and BERTScore frequently diverge from human quality assessments, creating a fundamental evaluation challenge. A breakthrough solution has emerged in the form of language model-based evaluation frameworks. Cortes et al. [21] found that carefully designed prompting strategies with advanced models like GPT-4 can achieve remarkable alignment with human judgment, particularly when assessing the thoroughness of responses. This paradigm shift suggests that AI systems themselves may offer the most effective tools for evaluating complex natural language generation tasks.
LLMs have become widely adopted as evaluators in benchmarks like AlpacaEval [23], MT-Bench [13], and Chatbot Arena [18], where models such as GPT-4 conduct pairwise preference assessments based on helpfulness, factuality, and engagement. Frameworks like G-Eval [43] further enable fine-grained scoring by prompting models to assess specific dimensions using structured rubrics. These LLM-based evaluations have shown stronger alignment with human judgments than traditional metrics, especially on long-form and knowledge-intensive tasks.
However, relying on a single model introduces concerns around opacity, bias, and instability, including self-preference and sensitivity to output variance. To address this, we implement a multi-model, multi-run evaluation protocols that improve interpretability, reduces bias, and yields more robust and reproducible assessments.
D.1.2 Leveraging an Interpretable Ensemble of Reasoning LLMs
To address the limitations of single-model, single-pass evaluation pipelines, we propose an interpretable and robust evaluation framework based on an ensemble of reasoning-capable LLMs: Deepseek-R1-Distilled [34], Qwen-3-32B [5], and Phi-4-Reasoning [6]. These models were selected not only for their demonstrated strength in multi-hop reasoning and long-form comprehension but also for their open accessibility, ensuring that our evaluation protocol is fully transparent and reproducible without dependence on proprietary APIs. Our evaluation protocol is also distinguished by its multi-run robustness. For each benchmark sample, we perform three independent inference runs of the candidate model to capture natural generation variability. Each of these outputs is then evaluated independently by the full ensemble of three judge models. This results in nine total evaluations per sample, allowing us to report aggregated metrics that reflect not only average performance but also stability and consistency across generations and evaluators. We further analyze cross-model agreement and judgment variance to ensure evaluation fidelity. This interpretable ensemble-based and multi-run evaluation represents a significant step forward in the use of LLMs as evaluators. It brings together the benefits of scale and automation while maintaining experimental rigor.
D.2 Reliability and Robustness of Multi-Judge Evaluation
To ensure the statistical robustness and reproducibility of our evaluation framework, we go beyond model-averaged scores and perform formal inter- and intra-judge reliability assessments. These analyses validate both the consistency across judges (inter-rater agreement) and stability within individual judges across multiple runs (intra-rater reliability), providing a more rigorous characterization of evaluation quality.
Inter-Judge Agreements: We assess agreement between our ensemble of LLM judges: Deepseek-R1-Distilled, Qwen-3-32B, and Phi-4-Reasoning, using two complementary statistical measures:
Fleiss’ Kappa: To measure categorical agreement across multiple judges, we binarize evaluation outcomes (e.g., correct vs. incorrect) and compute Fleiss’ [28], a widely used metric for evaluating agreement on nominal data among fixed raters. This quantifies how consistently the LLM judges classify outputs beyond what would be expected by chance.
Kendall’s W (Coefficient of Concordance: For tasks involving ordinal scoring or ranking of generations, we compute Kendall’s W [27], a non-parametric measure of rank correlation. This accounts for judges using different scoring scales by focusing on relative orderings. To accommodate tied ranks, we use the corrected-for-ties version of Kendall’s W. The coefficient ranges from 0 (no agreement) to 1 (perfect agreement), enabling us to quantify the degree of concordance in evaluative judgments.
The results of our inter-judge reliability analysis, shown in Figure 6, reveal consistently high agreement among the three LLM judges across a diverse set of evaluated models. Fleiss’ scores for the ID task (binary classification) generally fall within the 0.75–0.88 range, indicating “good” to “excellent” agreement by standard interpretation guidelines The bottom plot presents inter-judge agreement measured by Kendall’s W across four evaluation dimensions: accuracy, completeness, parsimony, and relevance, for 23 vision-language models. Overall, models exhibit moderate to strong agreement (W = 0.69–0.87), indicating consistent ranking behavior among the three LLM judges. Agreement is highest for accuracy. In contrast, parsimony scores demonstrate lower concordance (W = 0.64–0.73), indicating greater variability in how judges interpret brevity or conciseness. Qwen, Claude, and Gemma families show relatively stable agreement across all four dimensions, highlighting their reliability as evaluated agents. The relatively lower agreement for some models, such as LLaVA-based variants and InternVL3-2B, particularly on reasoning metrics, highlights instances where judge interpretations diverged, possibly due to varied output styles or ambiguous task completions.
Intra-Judge Reliability: In addition to evaluating agreement across different models, we assess intra-judge reliability by running each judge model three independent times on the same set of samples. This allows us to measure the stability of each model’s judgments under natural generation variability.
For this purpose, we compute the Intraclass Correlation Coefficient (ICC), which quantifies how consistently a single model scores the same item across multiple runs. ICC is particularly suited for this setting as it accounts for both within-subject and between-subject variability, providing a continuous-scale assessment of intra-rater consistency. Following established interpretative guidelines [20] we classify ICC values into bands (e.g., moderate, good, excellent) to report the strength of reliability for each judge.
To assess the consistency of our LLM-based judges, we conducted a three-run intra-rater reliability analysis using the Intraclass Correlation Coefficient (ICC2) shown in Figure 7, separately computed for binary ID judgments and ordinal reasoning assessments. The results indicate that both DeepSeek-R1 and Qwen3-32B exhibit strong intra-judge reliability. DeepSeek-R1 shows excellent agreement on ID assessments across almost all models, with ICC2 values typically in the 0.85–0.90 range. We also observe that LLaVA-based models tend to show more intra-model fluctuation, potentially due to less structured or variable outputs.
D.3 Diagnostic Parsimony
Diagnostic parsimony refers to the principled restraint in offering explanations, favoring the simplest account consistent with observed evidence while avoiding unnecessary speculation. [45] shows that humans have a robust bias toward simpler explanations, not just for cognitive ease but because simplicity aids understanding, memory, and decision-making. In agriculture, this principle is especially critical: farmers and gardeners often seek immediate, actionable advice under time-sensitive or resource-constrained conditions. Overly elaborate responses can obscure key insights, introduce confusion, or even lead to misapplied interventions. As in medicine and law, where expert communication must balance completeness with clarity, parsimony is a cornerstone of effective agronomic consultation. LLM/LVLM answers in such knowledge-intensive domains must therefore strive to be concise yet comprehensive, ensuring the user gets just enough information to act confidently. Studies show that human evaluators, particularly non-experts, tend to prefer shorter, to-the-point responses, even when some details are omitted [65].
However, many instruction-following LLMs/LVLMs often default to verbose outputs, including tangential or overly cautious elaborations. As illustrated in Figure 8, when responding to a user’s concern about bark defects on a tree, GPT-4.1 generates a detailed and structured diagnostic explanation covering wound physiology, tree tolerance, preventative care, and multi-step recommendations. Qwen 2 VL 72B, while more concise, still includes speculative possibilities (e.g., fungal infection, water stress) and a broad checklist of actions. In contrast, the human expert offers a brief but decisive interpretation, attributing the damage to non-serious activity and explicitly recommending no treatment. This behavior not only hinders user trust and satisfaction but also confounds traditional QA metrics like exact match (EM) and F1, which are poorly suited to judge informativeness under parsimony constraints. [9].
Appendix E Additional MMST Benchmark Results
E.1 MMST Benchmark Main Results
E.1.1 Standard ID Benchmark Results
Table 6 summarises identification accuracy (Acc, %) and reasoning accuracy (0–4 scale) for a diverse LVLM on the MMST Standard-ID benchmark, reporting scores under three automated judges—DeepSeek-R1-Distill, Qwen3-32B, and Phi-4-reasoning. The table reveals a persistent proprietary lead: GPT-4.1 achieves the highest scores across all judges, retaining a margin of roughly 13 pp in identification accuracy and 0.5 points in reasoning over the strongest open-source model, Qwen-2.5-VL-72B-Instruct.
| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||
| Acc (%) | Reasoning | Acc (%) | Reasoning | Acc (%) | Reasoning | |
| gpt-4.1 | 44.6 | 3.07 | 44.7 | 2.78 | 42.4 | 3.17 |
| gpt-4.1-mini | ||||||
| gpt-4o | ||||||
| gpt-4o-mini | ||||||
| claude-3-7-sonnet | ||||||
| claude-3-5-sonnet | ||||||
| claude-3-haiku | ||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||
| llava-v1.6-mistral-7b-hf | ||||||
| llava-onevision-qwen2-7b-ov-hf | 29 | 20 | 31 | 23 | 43 | 16 |
| Qwen2.5-VL-3B-Instruct | ||||||
| Qwen2.5-VL-7B-Instruct | 29 | 23 | 29 | 24 | 28 | 28 |
| Qwen2.5-VL-32B-Instruct | 16 | 33 | 9 | 28 | 16 | 32 |
| Qwen2.5-VL-72B-Instruct | 18 | 2 | 20 | 2 | 19 | 1 |
| gemma-3-4b-it | ||||||
| gemma-3-12b-it | 55 | 11 | 49 | 7 | 54 | 5 |
| gemma-3-27b-it | 20 | 10 | 21 | 12 | 15 | 15 |
| InternVL3-2B | ||||||
| InternVL3-8B | 22 | 10 | 37 | 7 | 39 | 5 |
| InternVL3-14B | 21 | 8 | 16 | 7 | 19 | 9 |
| InternVL3-38B | 36 | 9 | 39 | 11 | 31 | 12 |
| InternVL3-78B | 20 | 7 | 15 | 6 | 17 | 5 |
Models are color-coded by type: closed-source models in red/light red, open-source models in blue/light blue.
values indicate percentage improvements over the previous model size in the same family. Bold purple values highlight the best performance.
E.1.2 Standard MG Benchmark Results
Table 7 presents management-task performance on the MMST Standard-MG benchmark, reporting four rubric scores—Accuracy, Relevance, Completeness, and Parsimony (0–4 scale)—under the same trio of automated judges used for the ID setting. Consistent with the ID results, the proprietary GPT-4.1 model dominates most of metrics across all judges, outscoring the best open-source competitor by roughly absolute points in Accuracy and by – in the Relevance and Completeness.
| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||||||||
| Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | |
| gpt-4.1 | 3.17 | 3.59 | 3.39 | 3.27 | 3.59 | 3.03 | 3.27 | 3.62 | 3.24 | 3.14 | ||
| gpt-4.1-mini | ||||||||||||
| gpt-4o | 3.07 | |||||||||||
| gpt-4o-mini | ||||||||||||
| claude-3-7-sonnet | ||||||||||||
| claude-3-5-sonnet | ||||||||||||
| claude-3-haiku | ||||||||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
| llava-v1.6-mistral-7b-hf | ||||||||||||
| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
| Qwen2.5-VL-3B-Instruct | ||||||||||||
| Qwen2.5-VL-7B-Instruct | ||||||||||||
| Qwen2.5-VL-32B-Instruct | ||||||||||||
| Qwen2.5-VL-72B-Instruct | ||||||||||||
| gemma-3-4b-it | ||||||||||||
| gemma-3-12b-it | ||||||||||||
| gemma-3-27b-it | ||||||||||||
| InternVL3-2B | ||||||||||||
| InternVL3-8B | ||||||||||||
| InternVL3-14B | ||||||||||||
| InternVL3-38B | ||||||||||||
| InternVL3-78B | 2.97 | |||||||||||
Models are color-coded by type: closed-source models in red/light red, open-source models in blue/light blue. Scores represent performance on four key metrics (Accuracy / Relevance / Completeness / Parsimony) on a 0-4 scale. Bold purple values highlight the best performance on each metric within each benchmark.
E.1.3 Contextual MG Benchmark Results
| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||||||||
| Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | |
| gpt-4.1 | 3.21 | 3.61 | 3.40 | 3.28 | 3.59 | 3.01 | 3.29 | 3.61 | 3.24 | 3.14 | ||
| gpt-4.1-mini | ||||||||||||
| gpt-4o | 3.00 | |||||||||||
| gpt-4o-mini | ||||||||||||
| claude-3-7-sonnet | ||||||||||||
| claude-3-5-sonnet | ||||||||||||
| claude-3-haiku | ||||||||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
| llava-v1.6-mistral-7b-hf | ||||||||||||
| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
| Qwen2.5-VL-3B-Instruct | ||||||||||||
| Qwen2.5-VL-7B-Instruct | ||||||||||||
| Qwen2.5-VL-32B-Instruct | ||||||||||||
| Qwen2.5-VL-72B-Instruct | ||||||||||||
| gemma-3-4b-it | ||||||||||||
| gemma-3-12b-it | ||||||||||||
| gemma-3-27b-it | ||||||||||||
| InternVL3-2B | ||||||||||||
| InternVL3-8B | ||||||||||||
| InternVL3-14B | ||||||||||||
| InternVL3-38B | ||||||||||||
| InternVL3-78B | 3.05 | |||||||||||
Models are color-coded by type: closed-source models in red/light red, open-source models in blue/light blue.
Scores represent performance on four key metrics (Accuracy / Relevance / Completeness / Parsimony) on a 0-4 scale. Bold purple values denote the best score for each metric across all models.
E.2 Model Scaling Results
Increasing model scale consistently boosts both identification accuracy and reasoning quality for all three open-source LVLM families (See Figure 9).
E.3 With/Without Meta Data Results
Utility of metadata.
Tables 9 and 10 contrast model performance when the question is presented with versus without the user’s location and timestamp. For the Identification benchmark (Table 9), adding metadata yields only modest absolute gains ( pp ID%) and virtually no change in reasoning quality (). The strongest closed-source model (GPT-4.1) and the largest open-source Qwen-72B both extract a small benefit (+1.6 and +0.6 pp, respectively), whereas lighter counterparts (e.g., GPT-4.1-mini, Gemma-3B) see negligible or even negative shifts, suggesting that metadata can introduce noise unless the model possesses sufficient capacity to interpret it.
A similar pattern emerges on the Management benchmark (Table 10): absolute deltas across accuracy, relevance, completeness, and parsimony remain within , yet the direction of change is revealing. Large models (GPT-4.1, Gemma-27B, Qwen-72B) exhibit consistent, albeit small, improvements, most notably in relevance and completeness, while smaller models fluctuate or even decline. Overall, these findings indicate that spatiotemporal cues confer a measurable but limited advantage, and that leveraging them effectively remains contingent on model scale and training.
| Model | Image + Text Only | + Metadata | (Meta – No Meta) | |||
| ID% | Reason | ID% | Reason | ID% | Reason | |
| gpt-4.1 | 46.2 | 3.12 | ||||
| gpt-4.1-mini | 35.6 | 2.81 | ||||
| gemma-3-4b-it | 10.2 | 1.88 | ||||
| gemma-3-27b-it | 19.0 | 2.33 | ||||
| Qwen2.5-VL-3B-Instruct | 18.0 | 1.57 | ||||
| Qwen2.5-VL-32B-Instruct | 25.7 | 2.54 | ||||
| Qwen2.5-VL-72B-Instruct | 31.4 | 2.59 | ||||
| Model | Image + Text Only | + Metadata | (Meta – No Meta) | |||||||||
| Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | A | R | C | P | |
| gpt-4.1 | 3.21 | 3.62 | 3.41 | 2.96 | ||||||||
| gpt-4.1-mini | 2.90 | 3.41 | 3.05 | 2.94 | ||||||||
| gemma-3-4b-it | 2.32 | 2.83 | 2.62 | 2.18 | ||||||||
| gemma-3-27b-it | 2.81 | 3.28 | 3.19 | 2.55 | ||||||||
| Qwen2.5-VL-3B-Instruct | 2.13 | 2.44 | 1.99 | 2.00 | ||||||||
| Qwen2.5-VL-32B-Instruct | 2.83 | 3.25 | 3.14 | 2.54 | ||||||||
| Qwen2.5-VL-72B-Instruct | 2.71 | 3.12 | 2.83 | 2.56 | ||||||||
E.4 Fine-Turning Results
Fine-tuning setup.
All models are adapted on the MMST standard training dataset, which contains 17 532 single-turn consultations, each paired with up to three images. Given the limited corpus size, we employ parameter-efficient LoRA fine-tuning: a global batch size of 128, lora_alpha = 64, lora_dropout = 0.05, and bfloat16 precision. Optimisation uses AdamW with a cosine learning-rate schedule and a warm-up ratio of 0.03. Hardware resources: one NVIDIA H200 GPU suffices for the QwenVL-2.5-3 B and 7 B models, while the 32 B model is trained on two H200 cards—enabling the full eight-epoch run.
Effect of LoRA fine-tuning.
Figure 10 tracks identification accuracy and reasoning accuracy on seen vs. unseen entities, as Qwen2.5-VL models undergo progressively longer LoRA fine-tuning. For both the 32 B (left) and 7 B (right) variants, the bulk of the improvement is realised within the first four epochs: ID accuracy on seen entities rises from 32.9% to 37.6% for 32 B and from 27.7% to 34.8% for 7 B. Beyond epoch 4 the gains plateau or slightly regress, hinting at diminishing returns and possible over-fitting to the fine-tuning set. Reasoning accuracy follow a similar but more muted trend, increasing by at most 0.2–0.3 points before flattening.
The persistently low curves for unseen entities are unsurprising. Identification requires the model to emit an explicit entity name; if that name never appeared in the fine-tuning set, fine-tuning does not help.
E.5 Category-Wise Breakdown Results
[…]
E.5.1 Plant Identification Reults (MMST Standard)
| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||
| Acc (%) | Reasoning | Acc (%) | Reasoning | Acc (%) | Reasoning | |
| gpt-4.1 | 48.7 | 3.13 | 49.0 | 2.89 | 45.9 | 3.28 |
| gpt-4.1-mini | ||||||
| gpt-4o | ||||||
| gpt-4o-mini | ||||||
| claude-3-7-sonnet | ||||||
| claude-3-5-sonnet | ||||||
| claude-3-haiku | ||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||
| llava-v1.6-mistral-7b-hf | ||||||
| llava-onevision-qwen2-7b-ov-hf | ||||||
| Qwen2.5-VL-3B-Instruct | ||||||
| Qwen2.5-VL-7B-Instruct | ||||||
| Qwen2.5-VL-32B-Instruct | ||||||
| Qwen2.5-VL-72B-Instruct | ||||||
| gemma-3-4b-it | ||||||
| gemma-3-12b-it | ||||||
| gemma-3-27b-it | ||||||
| InternVL3-2B | ||||||
| InternVL3-8B | ||||||
| InternVL3-14B | ||||||
| InternVL3-38B | ||||||
| InternVL3-78B | ||||||
Models are color-coded by type: closed-source models in red/light red, open-source models in blue/light blue. Bold purple values denote the best performance in each column.
E.5.2 Insect and Pest Identification Results (MMST Standard)
| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||
| Acc (%) | Reasoning | Acc (%) | Reasoning | Acc (%) | Reasoning | |
| gpt-4.1 | 34.9 | 2.93 | 34.9 | 2.60 | 32.6 | 3.01 |
| gpt-4.1-mini | ||||||
| gpt-4o | ||||||
| gpt-4o-mini | ||||||
| claude-3-7-sonnet | ||||||
| claude-3-5-sonnet | ||||||
| claude-3-haiku | ||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||
| llava-v1.6-mistral-7b-hf | ||||||
| llava-onevision-qwen2-7b-ov-hf | ||||||
| Qwen2.5-VL-3B-Instruct | ||||||
| Qwen2.5-VL-7B-Instruct | ||||||
| Qwen2.5-VL-32B-Instruct | ||||||
| Qwen2.5-VL-72B-Instruct | ||||||
| gemma-3-4b-it | ||||||
| gemma-3-12b-it | ||||||
| gemma-3-27b-it | ||||||
| InternVL3-2B | ||||||
| InternVL3-8B | ||||||
| InternVL3-14B | ||||||
| InternVL3-38B | ||||||
| InternVL3-78B | ||||||
Models are color-coded by type: closed-source models in red/light red, open-source models in blue/light blue. Bold purple values denote the best performance in each column.
E.5.3 Plant Disease Identification Results (MMST Standard)
| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||
| Acc (%) | Reasoning | Acc (%) | Reasoning | Acc (%) | Reasoning | |
| gpt-4.1 | 45.7 | 3.06 | 44.6 | 2.62 | 46.0 | 2.97 |
| gpt-4.1-mini | ||||||
| gpt-4o | ||||||
| gpt-4o-mini | ||||||
| claude-3-7-sonnet | ||||||
| claude-3-5-sonnet | ||||||
| claude-3-haiku | ||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||
| llava-v1.6-mistral-7b-hf | ||||||
| llava-onevision-qwen2-7b-ov-hf | ||||||
| Qwen2.5-VL-3B-Instruct | ||||||
| Qwen2.5-VL-7B-Instruct | ||||||
| Qwen2.5-VL-32B-Instruct | ||||||
| Qwen2.5-VL-72B-Instruct | ||||||
| gemma-3-4b-it | ||||||
| gemma-3-12b-it | ||||||
| gemma-3-27b-it | ||||||
| InternVL3-2B | ||||||
| InternVL3-8B | ||||||
| InternVL3-14B | ||||||
| InternVL3-38B | ||||||
| InternVL3-78B | ||||||
Models are color-coded by type: closed-source models in red/light red, open-source models in blue/light blue. Bold purple values denote the best performance in each column.
E.5.4 Plant Disease Management Results (MMST Standard)
| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||||||||
| Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | |
| gpt-4.1 | 3.05 | 3.54 | 3.31 | 2.87 | 3.17 | 3.54 | 2.90 | 2.92 | 3.15 | 3.53 | 3.12 | 3.05 |
| gpt-4.1-mini | ||||||||||||
| gpt-4o | ||||||||||||
| gpt-4o-mini | ||||||||||||
| claude-3-7-sonnet | ||||||||||||
| claude-3-5-sonnet | ||||||||||||
| claude-3-haiku | ||||||||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
| llava-v1.6-mistral-7b-hf | ||||||||||||
| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
| Qwen2.5-VL-3B-Instruct | ||||||||||||
| Qwen2.5-VL-7B-Instruct | ||||||||||||
| Qwen2.5-VL-32B-Instruct | ||||||||||||
| Qwen2.5-VL-72B-Instruct | ||||||||||||
| gemma-3-4b-it | ||||||||||||
| gemma-3-12b-it | ||||||||||||
| gemma-3-27b-it | ||||||||||||
| InternVL3-2B | ||||||||||||
| InternVL3-8B | ||||||||||||
| InternVL3-14B | ||||||||||||
| InternVL3-38B | ||||||||||||
| InternVL3-78B | ||||||||||||
Models are color-coded by type: closed-source models in red/light red, open-source models in blue/light blue. Scores are given on a 0–4 scale for Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars). Bold purple numbers denote the best performance for each metric within a column block.
E.5.5 Insect and Pest Management Results (MMST Standard)
| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||||||||
| Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | |
| gpt-4.1 | 3.05 | 3.51 | 3.33 | 2.87 | 3.18 | 3.53 | 2.98 | 2.98 | 3.11 | 3.52 | 3.15 | 3.06 |
| gpt-4.1-mini | ||||||||||||
| gpt-4o | ||||||||||||
| gpt-4o-mini | ||||||||||||
| nova-pro | ||||||||||||
| claude-3-7-sonnet | ||||||||||||
| claude-3-5-sonnet | ||||||||||||
| claude-3-haiku | ||||||||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
| llava-v1.6-mistral-7b-hf | ||||||||||||
| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
| Qwen2.5-VL-3B-Instruct | ||||||||||||
| Qwen2.5-VL-7B-Instruct | ||||||||||||
| Qwen2.5-VL-32B-Instruct | ||||||||||||
| Qwen2.5-VL-72B-Instruct | ||||||||||||
| gemma-3-4b-it | ||||||||||||
| gemma-3-12b-it | ||||||||||||
| gemma-3-27b-it | ||||||||||||
| InternVL3-2B | ||||||||||||
| InternVL3-8B | ||||||||||||
| InternVL3-14B | ||||||||||||
| InternVL3-38B | ||||||||||||
| InternVL3-78B | ||||||||||||
Models are color-coded by type: closed-source (red) and open-source (blue). Scores range from 0-4 for Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars). Bold purple indicates the best score in each column block.
E.5.6 Plant Care and Gardening Guidance Results (MMST Standard)
| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||||||||
| Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | |
| gpt-4.1 | 3.38 | 3.73 | 3.55 | 3.01 | 3.45 | 3.73 | 3.17 | 3.12 | 3.52 | 3.80 | 3.43 | 3.24 |
| gpt-4.1-mini | ||||||||||||
| gpt-4o | ||||||||||||
| gpt-4o-mini | ||||||||||||
| nova-pro | ||||||||||||
| claude-3-7-sonnet | ||||||||||||
| claude-3-5-sonnet | ||||||||||||
| claude-3-haiku | ||||||||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
| llava-v1.6-mistral-7b-hf | ||||||||||||
| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
| Qwen2.5-VL-3B-Instruct | ||||||||||||
| Qwen2.5-VL-7B-Instruct | ||||||||||||
| Qwen2.5-VL-32B-Instruct | ||||||||||||
| Qwen2.5-VL-72B-Instruct | ||||||||||||
| gemma-3-4b-it | ||||||||||||
| gemma-3-12b-it | ||||||||||||
| gemma-3-27b-it | ||||||||||||
| InternVL3-2B | ||||||||||||
| InternVL3-8B | ||||||||||||
| InternVL3-14B | ||||||||||||
| InternVL3-38B | ||||||||||||
| InternVL3-78B | ||||||||||||
Models are color-coded by type: closed-source (red) and open-source (blue). Each metric is scored on a 0–4 scale: Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars). Bold purple values indicate the best performance within each column block.
E.5.7 Weeds/Invasive Plants Management Results (MMST Standard)
| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||||||||
| Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | |
| gpt-4.1 | 2.90 | 3.41 | 3.14 | 3.01 | 3.37 | 2.87 | 2.89 | 3.40 | 3.00 | |||
| gpt-4.1-mini | ||||||||||||
| gpt-4o | 2.95 | 3.17 | 3.11 | |||||||||
| gpt-4o-mini | ||||||||||||
| nova-pro | ||||||||||||
| claude-3-7-sonnet | ||||||||||||
| claude-3-5-sonnet | ||||||||||||
| claude-3-haiku | ||||||||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
| llava-v1.6-mistral-7b-hf | ||||||||||||
| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
| Qwen2.5-VL-3B-Instruct | ||||||||||||
| Qwen2.5-VL-7B-Instruct | ||||||||||||
| Qwen2.5-VL-32B-Instruct | ||||||||||||
| Qwen2.5-VL-72B-Instruct | ||||||||||||
| gemma-3-4b-it | ||||||||||||
| gemma-3-12b-it | ||||||||||||
| gemma-3-27b-it | ||||||||||||
| InternVL3-2B | ||||||||||||
| InternVL3-8B | ||||||||||||
| InternVL3-14B | ||||||||||||
| InternVL3-38B | ||||||||||||
| InternVL3-78B | ||||||||||||
Models are color-coded by type: closed-source (red) and open-source (blue). Scores range from 0–4 across Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars). Bold purple marks the highest score for each metric within a column block.
E.5.8 Plant Disease Management Results (MMST Contextual)
| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||||||||
| Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | |
| gpt-4.1 | 3.07 | 3.52 | 3.34 | 2.86 | 3.16 | 3.49 | 2.91 | 3.09 | 3.44 | 3.08 | 2.96 | |
| gpt-4.1-mini | ||||||||||||
| gpt-4o | 3.17 | |||||||||||
| gpt-4o-mini | ||||||||||||
| nova-pro | ||||||||||||
| claude-3-7-sonnet | ||||||||||||
| claude-3-5-sonnet | ||||||||||||
| claude-3-haiku | ||||||||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
| llava-v1.6-mistral-7b-hf | ||||||||||||
| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
| Qwen2.5-VL-3B-Instruct | ||||||||||||
| Qwen2.5-VL-7B-Instruct | ||||||||||||
| Qwen2.5-VL-32B-Instruct | ||||||||||||
| Qwen2.5-VL-72B-Instruct | ||||||||||||
| gemma-3-4b-it | ||||||||||||
| gemma-3-12b-it | ||||||||||||
| gemma-3-27b-it | ||||||||||||
| InternVL3-2B | ||||||||||||
| InternVL3-8B | ||||||||||||
| InternVL3-14B | ||||||||||||
| InternVL3-38B | ||||||||||||
| InternVL3-78B | ||||||||||||
Models are color-coded by type: closed-source (red) and open-source (blue). Scores range from 0–4 across Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars). Bold purple marks the highest score for each metric within a column block.
E.5.9 Insect and Pest Management Results (MMST Contextual)
| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||||||||
| Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | |
| gpt-4.1 | 3.14 | 3.59 | 3.32 | 2.99 | 3.25 | 3.61 | 2.99 | 3.06 | 3.22 | 3.60 | 3.20 | 3.21 |
| gpt-4.1-mini | ||||||||||||
| gpt-4o | ||||||||||||
| gpt-4o-mini | ||||||||||||
| nova-pro | ||||||||||||
| claude-3-7-sonnet | ||||||||||||
| claude-3-5-sonnet | ||||||||||||
| claude-3-haiku | ||||||||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
| llava-v1.6-mistral-7b-hf | ||||||||||||
| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
| Qwen2.5-VL-3B-Instruct | ||||||||||||
| Qwen2.5-VL-7B-Instruct | ||||||||||||
| Qwen2.5-VL-32B-Instruct | ||||||||||||
| Qwen2.5-VL-72B-Instruct | ||||||||||||
| gemma-3-4b-it | ||||||||||||
| gemma-3-12b-it | ||||||||||||
| gemma-3-27b-it | ||||||||||||
| InternVL3-2B | ||||||||||||
| InternVL3-8B | ||||||||||||
| InternVL3-14B | ||||||||||||
| InternVL3-38B | ||||||||||||
| InternVL3-78B | ||||||||||||
Models are color-coded by type: closed-source (red) and open-source (blue). Scores range from 0–4 for Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars). Bold purple highlights the best score for each metric within a column block.
E.5.10 Plant Care and Gardening Guidance Results (MMST Contextual)
| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||||||||
| Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | |
| gpt-4.1 | 3.40 | 3.73 | 3.54 | 3.46 | 3.72 | 3.16 | 3.53 | 3.80 | 3.42 | 3.27 | ||
| gpt-4.1-mini | ||||||||||||
| gpt-4o | 3.08 | |||||||||||
| gpt-4o-mini | ||||||||||||
| nova-pro | ||||||||||||
| claude-3-7-sonnet | ||||||||||||
| claude-3-5-sonnet | ||||||||||||
| claude-3-haiku | ||||||||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
| llava-v1.6-mistral-7b-hf | ||||||||||||
| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
| Qwen2.5-VL-3B-Instruct | ||||||||||||
| Qwen2.5-VL-7B-Instruct | ||||||||||||
| Qwen2.5-VL-32B-Instruct | ||||||||||||
| Qwen2.5-VL-72B-Instruct | ||||||||||||
| gemma-3-4b-it | ||||||||||||
| gemma-3-12b-it | ||||||||||||
| gemma-3-27b-it | ||||||||||||
| InternVL3-2B | ||||||||||||
| InternVL3-8B | ||||||||||||
| InternVL3-14B | ||||||||||||
| InternVL3-38B | ||||||||||||
| InternVL3-78B | 3.50 | |||||||||||
Models are color-coded by type: closed-source (red) and open-source (blue). Scores range from 0–4 for Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars). Bold purple highlights the best score for each metric within a column block.
E.5.11 Weeds/Invasive Plants Management Results (MMST Contextual)
| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||||||||
| Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | |
| gpt-4.1 | 3.03 | 3.53 | 3.27 | 2.92 | 3.12 | 3.46 | 2.92 | 3.03 | 3.02 | 3.45 | 3.07 | 3.20 |
| gpt-4.1-mini | ||||||||||||
| gpt-4o | ||||||||||||
| gpt-4o-mini | ||||||||||||
| nova-pro | ||||||||||||
| claude-3-7-sonnet | ||||||||||||
| claude-3-5-sonnet | ||||||||||||
| claude-3-haiku | ||||||||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
| llava-v1.6-mistral-7b-hf | ||||||||||||
| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
| Qwen2.5-VL-3B-Instruct | ||||||||||||
| Qwen2.5-VL-7B-Instruct | ||||||||||||
| Qwen2.5-VL-32B-Instruct | ||||||||||||
| Qwen2.5-VL-72B-Instruct | ||||||||||||
| gemma-3-4b-it | ||||||||||||
| gemma-3-12b-it | ||||||||||||
| gemma-3-27b-it | ||||||||||||
| InternVL3-2B | ||||||||||||
| InternVL3-8B | ||||||||||||
| InternVL3-14B | ||||||||||||
| InternVL3-38B | ||||||||||||
| InternVL3-78B | ||||||||||||
Models are color-coded by type: closed-source (red) and open-source (blue). Scores range from 0–4 for Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars). Bold purple marks the highest score for each metric within a column block.
Appendix F MIRAGE-MMMT
F.1 Benchmark Details
The MIRAGE-MMMT dataset as shown in Table 22, contains 861 multi-turn samples, each annotated with a high-level decision label—either Clarify (56.6%) or Respond (43.4%)—reflecting the expert’s intent in continuing the consultation. On average, each sample includes 2.11 images and spans 1.52 turns, capturing compact yet information-rich interactions.
| Overall Statistics | Total |
| Total Samples | 861 |
| Decision Distribution | |
| Clarify | 487 (56.6%) |
| Respond | 374 (43.4%) |
| Per-Sample Statistics | |
| Avg. Images per Sample | 2.11 |
| Avg. Turns per Sample | 1.52 |
| Word Count Statistics | |
| Avg. User-turn Words | 109.91 |
| Avg. Expert-turn Words | 80.57 |
| Distribution Statistics | |
| Max Images per Sample | 3 |
| Max Turns per Sample | 14 |
| Max User-turn Words | 1 488 |
| Max Expert-turn Words | 287 |
User and expert utterances are relatively verbose, with average lengths of 109.9 and 80.6 words respectively, and a maximum of 1,488 words in a user turn. Each sample includes up to 3 images and 14 turns, reflecting a wide range of complexity and interaction depth. These characteristics make the dataset well-suited for studying decision-making, goal inference, and clarification strategies in visually grounded, expert-guided dialogues.
F.2 Task Definition
MIRAGE-MT is a multimodal, multi-turn benchmark designed to evaluate conversational expert agents in a consultative decision-making setting. Given a multi-turn dialogue and associated image(s) as shown in Figure 11, the agent must decide whether to ask a clarification question or provide a helpful response, and then generate the corresponding utterance.
F.2.1 Input
Let a multi-turn dialogue context be represented as a sequence:
where:
denotes the speaker
is the corresponding utterance
Each dialogue is also associated with a set of image inputs:
which may provide visual context necessary for interpretation (e.g., pest damage, plant structure).
F.2.2 Output
The model must jointly predict:
1.
A decision
2.
A corresponding utterance , where:
If , then is a clarification question
If , then is an expert answer
F.2.3 Goal Inference and Decision Policy
Let denote the user’s underlying goal (e.g., identifying a plant disease, choosing a planting strategy). The model must infer and a goal-state representation:
The model learns a policy:
and must select the appropriate action:
The generation should then follow:
F.3 Evaluation Criteria
Predictions are evaluated against:
Gold revealed fact , obtained from the masked user utterance after expert’s turn
Gold goal state , obtained from the source dialogue
LLM-as-a-Judge ratings:
Decision Accuracy
Goal Relevance
F.4 Data Curation Details
Each task sample consists of the dialogue context, referenced images, and metadata such as source dialogue ID and topic. To ensure data safety, we perform automated PII sanitization, replacing all named entities with randomized placeholders while preserving domain relevance. URLs and institutional references are retained when necessary for contextual fidelity. To ensure data quality and task validity, we conduct manual human review on a representative subset of the generated examples. Expert annotators assess the correctness of the decision label, coherence of the generated question or response, and alignment with the revealed user intent. Feedback from this process is used to refine prompt instructions and filter any low-quality generations. Our modular pipeline supports deterministic regeneration of the dataset via fixed seeds and indexing, enabling reproducible experimentation and future extensibility to other domains.
Release Protocol for MIRAGE-MMMT: In designing our dataset release, we follow established best practices from recent benchmarks such as MMLU [35], and BIG-Bench [32, 33], which emphasize the importance of separating training data and test targets to prevent leakage and ensure reliable model evaluation. We adopt a protocol that maximizes transparency, reproducibility, and community usability, while preserving the integrity of the held-out test set. We publicly release:
Full training and validation task datasets, generated from processed source conversations
Corresponding dialogue context, goal annotations, image references, and model-generated outputs
Task generation scripts, PII scrubbing utilities, and evaluation tools.
To ensure the credibility and integrity of the test set, we do not plan to release the source dialogues or revealed facts used to construct it. Instead, we provide only the test input (dialogue context and image references). This ensures that models are evaluated blind to the gold output, preventing overfitting or prompt leakage. Evaluation of model predictions on the test set can be conducted either via our LLM-based judge or via human assessment.
F.5 Additional MIRAGE-MMMT Results
| Input Variant | Decision Acc. | F1 (Macro) | Level | ||
| Dialogue only | 69.79% | 0.70 | Realistic | ||
| Dialogue + Goal | 71.34% | 1.55% | 0.71 | 0.01 | Semi-Privileged |
| Dialogue + GoalState | 89.27% | 19.48% | 0.89 | 0.19 | Oracle |
Appendix G Prompts
G.1 Evaluation Prompts for MIRAGE-MMST
Figure 13 presents the inference prompt used for the MIRAGE-MMST Identification Task. user_query refers to the original user question. This task evaluates both the model’s identification accuracy and reasoning quality, requiring it to generate a clear reasoning chain followed by a final answer. The prompt enforces a standardized output format to facilitate consistent and automatic evaluation. In contrast, for the management task, we impose no format constraints—models are simply given the user question along with the associated images during inference.
Figure 14 presents the evaluation prompt used for the MIRAGE-MMST Identification Task. Here, entity_type denotes the category of the entity—plant, disease, or insect/pest. user_query is the original user question, while expert_answer contains the expert’s full response. The field entity_name captures the specific entity mentioned by the expert, with its corresponding scientific name stored in entity_scientific_name. The list of entity_common_names comprises commonly used names for that entity, collected through external search. Finally, model_response refers to the generated answer being evaluated.
Figure 15 presents the evaluation prompt used for the MIRAGE-MMST Management Task. Here, user_query is the original user question, while expert_answer contains the expert’s full response. The field model_response refers to the generated answer being evaluated.
G.2 Evaluation Prompts for MIRAGE-MMMT
Appendix H Case Study
[…]
H.1 Category-Wise Cases
H.1.1 Plant Identification (MMST Standard)
H.1.2 Insect and Pest Identification (MMST Standard)
H.1.3 Plant Disease Identification (MMST Standard)
H.1.4 Plant Disease Management (MMST Standard)
H.1.5 Insect and Pest Management (MMST Standard)
H.1.6 Plant Care and Gardening Guidance (MMST Standard)
H.1.7 Weeds/Invasive Plants Management (MMST Standard)
H.1.8 Plant Disease Management (MMST Contextual)
H.1.9 Insect and Pest Management (MMST Contextual)
H.1.10 Plant Care and Gardening Guidance (MMST Contextual)
H.1.11 Weeds/Invasive Plants Management (MMST Contextual)
H.1.12 Clarify (MMMT)
H.1.13 Respond (MMMT)
H.2 Examples of Reasoning LLM as a Judge
H.2.1 MMST Identification Task
H.2.2 MMST Management Task
H.2.3 MMMT Task

### v2

@@
 1 Introduction
-Advances in large vision–language models (LVLMs) have significantly improved AI’s ability to interpret images and generate natural-language responses.
-However, existing benchmarks predominantly focus on short-form visual question answering [68, 30, 12, 51], captioning [53, 50], or grounded generation under constrained contexts [29].
+Advances in large vision–language models (LVLMs) have significantly improved AI’s ability to interpret images and generate natural‐language responses.
+However, existing benchmarks predominantly focus on short-form visual question answering [yue2024mmmu, gauba2025agmmu, antol2015vqa, schwenk2022okvqa, zhang2024b], captioning [singla2024pixels, schuhmann2021laion], or grounded generation under constrained contexts [garcia2022improving].
 These tasks fall short of capturing the interactive, decision-centric nature of real-world expert consultations, where users often present open-ended, ambiguous, and visually grounded queries.
-In knowledge-intensive domains such as agriculture, medicine, and engineering, expert consultations inherently span multiple modalities [49] and success hinges not just on perception or language fluency, but on the ability to reason causally, handle missing context, and make interaction-level decisions.
+In knowledge-intensive domains such as agriculture, medicine, and engineering, expert consultations inherently span multiple modalities [salvi2024multi] and success hinges not just on perception or language fluency, but on the ability to reason causally, handle missing context, and make interaction-level decisions.
 This complex interplay between multimodal understanding and contextual reasoning in professional settings represents a significant gap in the current LVLM evaluation frameworks.
-LLMs and LVLMs are increasingly used in domains like healthcare, law, and plant care [57, 59, 38, 71, 17, 48, 63], yet current benchmarks underrepresent knowledge-intensive scenarios involving ambiguous, multimodal queries [30].
-Agriculture illustrates this gap: farmers often seek image-based, expert-level guidance during critical events, but inaccurate or unsupported model outputs [69, 40, 67] can lead to serious consequences.
+LLMs and LVLMs are increasingly used in domains like healthcare, law, and plant care [tomar2024ecosage, wang2024agri, li2023llava, zhang2023huatuogpt, chen2024towards, ren2024healthcare, wu2024knowledge], yet current benchmarks underrepresent knowledge-intensive scenarios involving ambiguous, multimodal queries [gauba2025agmmu].
+Agriculture illustrates this gap: farmers often seek image-based, expert-level guidance during critical events, but inaccurate or unsupported model outputs [zhai2023halle, liu2024survey, yin2024woodpecker] can lead to serious consequences.
 This highlights the need for rigorous evaluation frameworks that assess VLMs in more realistic settings.
@@
 2.1 Overview of MIRAGE
-MIRAGE comprises two components: MMST, with over 29,000 single-turn expert-user interactions, and MMMT, a multi-turn corpus of 6,306 dialogues ( 12,000 turns) requiring clarify-or-respond decisions.
+MIRAGE comprises two components: MMST, with a total corpus of over 29,000 single-turn expert-user interactions, and MMMT, a multi-turn corpus of 6,306 dialogues ( 12,000 turns) requiring clarify-or-respond decisions.
+Figure 1 provides a visual overview of the evaluation benchmarks derived from this corpus.
 Tasks span identification, diagnostic reasoning, and issue management guidance across seven agronomic categories (Table 4).
-MMST includes Standard and Contextual subsets (Section 3, Appendix C), and a fine-grained evaluation framework assesses both accuracy and utility.
-With over 7,600 biological entities, including thousands of unseen species in the evaluation set, MIRAGE introduces a challenging open-world generalization setting for LVLMs.
+The MMST evaluation benchmark includes Standard (8,184 interactions) and Contextual (3,934 interactions) subsets (Section 3, Appendix C), while the MMMT benchmark contains 861 dialogues.
+A fine-grained evaluation framework assesses both accuracy and utility.
+With over 7,600 biological entities in the full corpus (6,856 in the benchmark set), including thousands of unseen species in the evaluation set, MIRAGE introduces a challenging open-world generalization setting for LVLMs.
 Refer Appendix C for more details.
 | Dataset | Type | Multimodal | Training Set | Expert Domain | Factuality | Expert Authored | Multi Turn |
-| iNat21 [58] | CLS | ✗ | ✓ | ✓ | – | – | – |
-| TreeOfLife [54] | CLS | ✓ | ✓ | ✓ | – | – | – |
-| SimpleQA [61] | OEQ | ✗ | ✗ | ✗ | ✓ | – | – |
-| MMMU [68] | MCQ | ✓ | ✗ | ✓ | ✗ | ✗ | ✗ |
-| AgMMU [30] | MCQ+OEQ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ |
-| CROP [70] | LFQ/Conv | ✗ | ✓ | ✓ | ✓ | ✓ | ✓ |
+| iNat21 [van2018inaturalist] | CLS | ✗ | ✓ | ✓ | – | – | – |
+| TreeOfLife [stevens2024bioclip] | CLS | ✓ | ✓ | ✓ | – | – | – |
+| SimpleQA [wei2024benchmarking] | OEQ | ✗ | ✗ | ✗ | ✓ | – | – |
+| MMMU [yue2024mmmu] | MCQ | ✓ | ✗ | ✓ | ✗ | ✗ | ✗ |
+| AgMMU [gauba2025agmmu] | MCQ+OEQ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ |
+| CROP [zhang2024empowering] | LFQ/Conv | ✗ | ✓ | ✓ | ✓ | ✓ | ✓ |
 | MIRAGE | LFQ/Conv | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
@@
 MIRAGE builds on a growing body of benchmarks in agriculture and multimodal reasoning, yet fills a critical gap left by prior efforts.
-Datasets like MMMU [68] focus on generalist multimodal reasoning but rely on constrained multiple-choice formats.
-AgMMU [30] targets agriculture, yet uses synthetic short-form MCQs that are not representative of real expert-user interactions.
-Domain-specific datasets such as TreeOfLife [54] and iNat21 [58] emphasize fine-grained image classification but operate in closed-world, non-interactive settings.
-The CROP benchmark [70] introduces multi-turn crop science QA, but is limited to two crops, lacks visual input, and is fully text-based.
+Datasets like MMMU [yue2024mmmu] focus on generalist multimodal reasoning but rely on constrained multiple-choice formats.
+AgMMU [gauba2025agmmu] targets agriculture, yet uses synthetic short-form MCQs that are not representative of real expert-user interactions.
+Domain-specific datasets such as TreeOfLife [stevens2024bioclip] and iNat21 [van2018inaturalist] emphasize fine-grained image classification but operate in closed-world, non-interactive settings.
+The CROP benchmark [zhang2024empowering] introduces multi-turn crop science QA, but is limited to two crops, lacks visual input, and is fully text-based.
 In contrast, MIRAGE is the only benchmark constructed from large-scale, real-world agricultural consultations and supports both multimodal single-turn and multi-turn tasks grounded in naturally underspecified user queries.
@@
 MIRAGE-MMST is a benchmark designed to assess expert-level, single-turn reasoning in multimodal agricultural consultations.
-The task setup is similar to a Long-form VQA task [37].
+The task setup is similar to a Long-form VQA task [huh2024long].
 Each instance consists of a natural language question paired with one or more user-provided images and associated metadata (e.g., timestamp, location).
@@
 To support varying levels of difficulty and contextual grounding, MIRAGE-MMST is divided into two subsets: a Standard subset, consisting of self-contained questions that can be answered using only the provided text and image, and a Contextual subset, where successful interpretation depends on implicit information such as time, location, or agricultural context not present in the input.
-The standard subset evaluates a model’s ability to identify entities, infer causal relationships, and generate explanatory or recommendation-based responses.
+The Standard subset further includes two task types—MMST-ID (Identification), which focuses on visual reasoning and recognizing entities, and MMST-MG (Management), which involves reasoning and generating recommendation-based responses.
 In contrast, the Contextual subset contains queries with latent information gaps and elliptical language, requiring models to reconstruct missing context using external priors.
@@
 To ensure data safety, we perform automated PII sanitization while preserving domain relevance.
-(see Appendix F.2 for more details)
+(see Appendix G.2 for more details)
 5 Experiments
@@
 5.1 Model Selection
-The LVLMs we evaluated can be divided into two groups: 1.) Proprietary Models: this group comprises SOTA GPT models (4.1, 4.1-mini, 4o, 4o-mini) [7], Claude-3 (3.5v2 Sonnet, 3.7 Sonnet, 3 Haiku) [11] available through API service.
-2.) Open-Source Models: this group includes LLaMa-4 Scout-17B [44], LLaVa (7B mistral, 7B qwen2) [41], Qwen 2.5 VL models (3B, 7B, 32B, 72B) [5], Gemma 3 models (4b, 12, 27b) [56] and Intern VL3 (2B, 8B, 14B, 38B, 72B) [73].
+The LVLMs we evaluated can be divided into two groups: 1.) Proprietary Models: this group comprises SOTA GPT models (4.1, 4.1-mini, 4o, 4o-mini) [achiam2023gpt], Claude-3 (3.5v2 Sonnet, 3.7 Sonnet, 3 Haiku) [anthropic2024claude] available through API service.
+2.) Open-Source Models: this group includes LLaMa-4 Scout-17B [metaai2025llama4], LLaVa (7B mistral, 7B qwen2) [liu2023visual], Qwen 2.5 VL models (3B, 7B, 32B, 72B) [githubQwen3Qwen3_Technical_ReportpdfMain], Gemma 3 models (4b, 12, 27b) [team2025gemma] and Intern VL3 (2B, 8B, 14B, 38B, 72B) [zhu2025internvl3].
 3.) Finetuned Models We perform multi-image multimodal fine-tuning of the Qwen 2.5 VL models on the MMST dataset.
@@
 To facilitate consistent, interpretable, and reproducible evaluation of model outputs, we implement a structured LLM-as-a-Judge protocol.
-Rather than relying on single-model, single-pass assessments, we leverage an ensemble of reasoning-capable language models: DeepSeek-R1-Distilled [34], Qwen3-32B [5], and Phi-4-Reasoning [6], as judges.
+Rather than relying on single-model, single-pass assessments, we leverage an ensemble of reasoning-capable language models: DeepSeek-R1-Distilled [guo2025deepseek], Qwen3-32B [githubQwen3Qwen3_Technical_ReportpdfMain], and Phi-4-Reasoning [abdin2025phi], as judges.
 Each candidate’s response is evaluated across three independent generations, with every generation scored by the full ensemble, resulting in nine total evaluations per sample.
@@
 We aggregate these scores in a 2:1:1:1 ratio.
-See Appendix D.1.2 for more details.
+See Appendix D for more details.
 5.4 Performance Comparison
@@
 | Model | MMST-ID Task | MMST-MG Task | |||||
-| Acc (%) | Reason | Acc | Rel | Comp | Pars | W-Sum | |
+| Acc (%) | Reason | Acc | Rel | Comp | Pars | W‑Sum | |
 | gpt-4.1 | 43.9 | 3.01 | 3.24 | 3.60 | 3.22 | 3.01 | 0.82 |
@@
 | InternVL3-78B | |||||||
-Scores are averaged over the three judge models. Closed-source rows are red/light-red, open-source rows blue/light-blue. Bold purple numbers mark the overall best for each metric.
+Scores are averaged over the three judge models. Closed‑source rows are red/light‑red, open‑source rows blue/light‑blue. Bold purple numbers mark the overall best for each metric.
 Performance improves steadily with model scale within families, especially for open models as seen in Figure 9.
@@
 Model Finetuning: We utilize the MMST training dataset to fine-tune Qwen 2.5 VL models (3B, 7B, 32B) and evaluate performance on both MMST-ID & MMST-MG tasks.
-Figure 1 illustrates the effects of LoRA-based fine-tuning on the Qwen2.5-VL-3B model across five checkpoints.
+Figure 2 illustrates the effects of LoRA-based fine-tuning on Qwen2.5-VL-3B model across four checkpoints vs the base model (base-Instruct).
 Our results indicate consistent improvements in seen entity accuracy on MMST-ID after fine-tuning, rising from 22.3% (base-Instruct) to a peak of 28.4% (epoch 6).
@@
 5.6 Performance Comparison
-We evaluate the model’s high-level decision and its alignment with user goals in three complementary ways.
+We evaluate the model’s high‐level decision and its alignment with user goals in three complementary ways.
 First, Decision Accuracy captures how often the model’s chosen action, whether to ask a clarification question or to respond directly, matches the gold annotation; To ensure that when the model generates an utterance, it is doing so meaningfully, we measure Clarify Relevance: for every turn labeled <Clarify> or Respond Relevance for all turns labeled <Respond> using the LLM judges.
@@
 | Qwen 72B | |||||||||
-We conduct experiments under Zero-Shot and COT [60] prompt settings on MMMT shown in Table 3.
+We conduct experiments under Zero-Shot and COT [wei2022chain] prompt settings on MMMT shown in Table 3.
 In the zero-shot setting, GPT-4o achieves the highest Decision Accuracy (62.98%) and leads in both Clarify Goal Relevance (70.75%) and Respond Goal Relevance (77.07%), showing strong overall reasoning and generation alignment.
@@
 Improving performance on unseen entities will require better open-world generalization strategies, such as retrieval-augmented generation, compositional reasoning, or domain-adaptive pretraining.
-Limited gains from metadata highlight the need for explicit context modeling—including structured encodings of time, location, and integration with external knowledge bases capturing seasonal and biological priors.
+Limited gains from metadata highlight the need for explicit context modeling, including structured encodings of time, location, and integration with external knowledge bases capturing seasonal and biological priors.
 Persistent challenges in clarify-or-respond decision-making point to a need for models that can infer latent user goals and reason about missing information.
@@
 MIRAGE is publicly released to support the development of vision-language systems that go beyond narrow question answering, toward models capable of engaging in natural interactions that involve ambiguity, a need for clarification, visual understanding, and decision-making when critical context is implicit or missing, addressing common challenges in real-world expert consultations.
-7 Limitations
-While MIRAGE provides a comprehensive evaluation framework for multimodal expert-level reasoning within agricultural consultations, several inherent limitations remain.
-First, the scope of the benchmark primarily reflects smallholder and backyard gardening scenarios, sourced largely from AskExtension’s database.
-Consequently, it does not fully encompass large-scale industrial agricultural practices or the more complex logistical, economic, and environmental issues prevalent in commercial farming operations.
-Additionally, agriculture is an extensive and multidisciplinary domain, covering areas such as soil science, agricultural engineering, economics, environmental sustainability, and policy implications.
-The current dataset captures only a subset of these dimensions.
-In the multi-turn scenario (MMMT), although designed to assess models’ ability to reason over dialogue context, the task does not yet simulate dynamic, truly interactive exchanges between models and users.
-The evaluations focus primarily on the models’ capacity to interpret and infer latent contextual information and to determine appropriate conversational actions based on partial observability.
-Consequently, MIRAGE does not fully measure interactive conversational reasoning, such as handling conversational repairs or real-time user feedback.
+7 Limitations & Future Work
+While MIRAGE provides a comprehensive evaluation framework for multimodal expert-level reasoning within agricultural consultations, several opportunities for future improvement remain.
+Currently, the benchmark primarily reflects smallholder and backyard gardening scenarios, largely derived from the AskExtension database.
+In future work, we aim to expand its scope to better represent large-scale industrial agricultural practices and the broader logistical, economic, and environmental challenges characteristic of commercial farming operations.
+Moreover, agriculture is inherently multidisciplinary, encompassing soil science, agricultural engineering, economics, environmental sustainability, and policy considerations.
+Future iterations of MIRAGE will extend coverage across these dimensions to better capture the full breadth of agricultural reasoning.
+In the multi-turn scenario (MMMT), our present design assesses models’ ability to reason over dialogue context, but it does not yet fully simulate dynamic, interactive exchanges between models and users.
+To address this, future versions will incorporate agentic capabilities that enable models to leverage time- and location-based context, as well as respond adaptively to evolving user input.
+Such extensions will allow MIRAGE to better evaluate interactive conversational reasoning, including conversational repairs, user feedback integration, and contextually grounded decision-making.
 8 Impact Statement
@@
 National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
+MIRAGE: A Benchmark for Multimodal Information-Seeking and Reasoning in Agricultural Expert-Guided Conversations
 Supplementary Material
 […]
-[…]
 Appendix A Data Source
-The MIRAGE benchmark is constructed from a large-scale archive of real-world agricultural consultations obtained from Ask Extension [25], a U.S.
+The MIRAGE benchmark is constructed from a large-scale archive of real-world agricultural consultations obtained from Ask Extension [extension_ask_extension], a U.S.
 national digital platform maintained by the Extension Foundation.
@@
 Each entry captures a real question from a user, along with the corresponding expert response, and may include user-uploaded images, time of submission, and geographic metadata.
-As seen in Figure 3, Our filtered AskExtension dialogues are strongly multimodal and reference-rich.
+As seen in Figure 4, Our filtered AskExtension dialogues are strongly multimodal and reference‐rich.
 Users typically include one to three images in their questions, about 35% of turns have a single image, 26% include two, and 39% include three.
@@
 Multimodal Large Language Models: Recent advances in multimodal large language models (LLMs) have markedly expanded vision–language reasoning capabilities.
-Proprietary models such as GPT-4 [8], Claude 3 Sonnet [10], and Google’s Gemini [55] demonstrate strong capabilities in unifying visual and textual modalities, achieving notable success across diverse multimodal benchmarks.
-Concurrently, open-source models—including Qwen-VL 2.5 [14], Gemma 3 [56], and InternVL-3 [73]—have narrowed the performance gap while remaining publicly accessible.
-Although these models excel on general-domain benchmarks, they underperform in agriculture: they lack fine-grained visual expertise, agronomic terminology, and the ability to reason about rare biological entities and management practices.
-MIRAGE is designed to expose these weaknesses by providing domain-specific, multimodal tasks that require expert-level diagnosis and advice.
-Multimodal Agricultural Benchmarks: Agricultural benchmarks [19, 26, 31] have progressed from controlled, image-only datasets to more realistic vision–language corpora.
-Early resources such as PlantVillage [36] and PlantDoc [52] emphasise leaf-level disease classification under laboratory and in-field conditions, respectively.
-The AgMMU [30] benchmark augments field images with farmer–expert dialogues, yet its evaluation is dominated by short, synthetic multiple-choice questions that limit open-ended reasoning.
-The recent Crop Disease Domain Multimodal (CDDM) [42] corpus contributes 137k crop–disease images paired with 1 M single-turn QA pairs, but it focuses narrowly on disease identification and management.
-In contrast, MIRAGE is derived from large-scale, real-world consultations covering more than 7000 biological entities; it offers both single-turn and multi-turn tasks, explicitly models clarify-or-respond decisions, and scores long-form answers along multiple dimensions of reasoning quality, thereby providing a far more ecologically valid and challenging test bed.
-Long-Form Question Answering Evaluation: Lexical-overlap metrics such as ROUGE [39], BLEU [46], and BERTScore [72] are ill-suited to evaluating open-ended, knowledge-intensive answers, as they correlate poorly with human judgements.
-Recent work therefore adopts LLM-as-Judge paradigms, exemplified by AlpacaEval [22], MT-Bench [13], and G-Eval [43], in which powerful LLMs are prompted to assess responses along axes such as factuality, coherence, and completeness.
-While these methods better align with human preferences, single-model evaluations remain vulnerable to bias and opacity.
-MIRAGE employs an interpretable ensemble of reasoning-focused LLM judges DeepSeek-R1-Distilled [34], Qwen 3-32B [5], and Phi-4-Reasoning [6] and by conducting three independent generation–evaluation passes per sample.
-This multi-model, multi-run protocol enhances robustness, enables variance analysis, and provides publicly inspectable rationales, yielding transparent and reproducible long-form QA evaluation.
-State Tracking in Dialogue Systems: State tracking has been a central component in task-oriented dialogue systems [16], where it plays a critical role in maintaining a representation of user intent and progress toward the user’s goal throughout the interaction.
+Proprietary models such as GPT-4 [gpt-4], Claude 3 Sonnet [claude-sonnet], and Google’s Gemini [team2023gemini] demonstrate strong capabilities in unifying visual and textual modalities, achieving notable success across diverse multimodal benchmarks.
+Concurrently, open‑source models—including Qwen‑VL 2.5 [bai2025qwen2], Gemma 3 [team2025gemma], and InternVL‑3 [zhu2025internvl3] have narrowed the performance gap while remaining publicly accessible.
+Although these models excel on general‑domain benchmarks, they underperform in agriculture: they lack fine‑grained visual expertise, agronomic terminology, and the ability to reason about rare biological entities and management practices.
+MIRAGE is designed to expose these weaknesses by providing domain‑specific, multimodal tasks that require expert‑level diagnosis and advice.
+Multimodal Agricultural Benchmarks: Agricultural benchmarks [agri_bench1, agri_bench2, agri_bench3] have progressed from controlled, image‑only datasets to more realistic vision–language corpora.
+Early resources such as PlantVillage [PlantVillage] and PlantDoc [singh2020plantdoc] emphasise leaf‑level disease classification under laboratory and in‑field conditions, respectively.
+The AgMMU [gauba2025agmmu] benchmark augments field images with farmer–expert dialogues, yet its evaluation is dominated by short, synthetic multiple‑choice questions that limit open‑ended reasoning.
+The recent Crop Disease Domain Multimodal (CDDM) [CDDM] corpus contributes 137k crop–disease images paired with 1 M single‑turn QA pairs, but it focuses narrowly on disease identification and management.
+In contrast, MIRAGE is derived from large‑scale, real‑world consultations covering more than 7000 biological entities; it offers both single‑turn and multi‑turn tasks, explicitly models clarify‑or‑respond decisions, and scores long‑form answers along multiple dimensions of reasoning quality, thereby providing a far more ecologically valid and challenging test bed.
+Long‑Form Question Answering Evaluation: Lexical‑overlap metrics such as ROUGE [rouge], BLEU [bleu], and BERTScore [zhang2019bertscore] are ill‑suited to evaluating open‑ended, knowledge‑intensive answers, as they correlate poorly with human judgements.
+Recent work therefore adopts LLM‑as‑Judge paradigms, exemplified by AlpacaEval [alpacaeval], MT‑Bench [bai2024mt], and G‑Eval [liu2023g], in which powerful LLMs are prompted to assess responses along axes such as factuality, coherence, and completeness.
+While these methods better align with human preferences, single‑model evaluations remain vulnerable to bias and opacity.
+MIRAGE employs an interpretable ensemble of reasoning‑focused LLM judges DeepSeek‑R1‑Distilled [guo2025deepseek], Qwen 3‑32B [githubQwen3Qwen3_Technical_ReportpdfMain], and Phi‑4‑Reasoning [abdin2025phi] and by conducting three independent generation–evaluation passes per sample.
+This multi‑model, multi‑run protocol enhances robustness, enables variance analysis, and provides publicly inspectable rationales, yielding transparent and reproducible long‑form QA evaluation.
+State Tracking in Dialogue Systems: State tracking has been a central component in task-oriented dialogue systems [chen2017survey, liu2024convbench], where it plays a critical role in maintaining a representation of user intent and progress toward the user’s goal throughout the interaction.
 Dialogue state tracking (DST) typically involves predicting a belief state, a structured representation of slot-value pairs reflecting the current user intent at every turn.
 The widespread adoption of state tracking in dialogue systems has enabled robust multi-turn interactions, improved task success rates, and enhanced user satisfaction.
-Benchmarks like MultiWOZ [15, 24, 66] and Schema-Guided Dialogue (SGD) [47] have been instrumental in shaping the field by introducing multi-domain, open-schema challenges that pushed models to become more generalizable and scalable.
+Benchmarks like MultiWOZ [budzianowski2018multiwoz, eric2019multiwoz, ye2021multiwoz] and Schema-Guided Dialogue (SGD) [rastogi2020towards] have been instrumental in shaping the field by introducing multi-domain, open-schema challenges that pushed models to become more generalizable and scalable.
 These successes in dialogue suggest that structured, continuously updated representations of task progress can be highly beneficial in other interactive or sequential decision-making settings.
@@
 | Total Images | 15 069 | 8 069 | 33 120 |
-| Per-Sample Statistics | |||
+| Per‑Sample Statistics | |||
 | Avg. Question Words | 69.57 | 80.94 | 67.53 |
@@
 Rather than standardizing to a single canonical form, our objective was to compile an exhaustive synonymy for each entity, enabling robust matching regardless of the nomenclatural variant used.
-The system queries multiple authoritative taxonomic databases via their APIs in a hierarchical approach, beginning with the Global Biodiversity Information Facility (GBIF) [2] and extending to iNaturalist [58] Encyclopedia of Life (EOL) [1], Integrated Taxonomic Information System (ITIS) [3], Wikipedia [62], and NCBI Taxonomy [4].
+The system queries multiple authoritative taxonomic databases via their APIs in a hierarchical approach, beginning with the Global Biodiversity Information Facility (GBIF) [GBIF] and extending to iNaturalist [van2018inaturalist] Encyclopedia of Life (EOL) [EOL], Integrated Taxonomic Information System (ITIS) [ITISgov], Wikipedia [Wikipedia], and NCBI Taxonomy [NIHTaxonomy].
 Queries are enhanced with category-specific context (e.g., kingdom Plantae for botanical entries) to improve retrieval accuracy.
@@
 Research by Xu et al.
-[64] reveals that established approaches like ROUGE and BERTScore frequently diverge from human quality assessments, creating a fundamental evaluation challenge.
+[criticalxu2023critical] reveals that established approaches like ROUGE and BERTScore frequently diverge from human quality assessments, creating a fundamental evaluation challenge.
 A breakthrough solution has emerged in the form of language model-based evaluation frameworks.
 Cortes et al.
-[21] found that carefully designed prompting strategies with advanced models like GPT-4 can achieve remarkable alignment with human judgment, particularly when assessing the thoroughness of responses.
+[beyondcortes2024beyond] found that carefully designed prompting strategies with advanced models like GPT-4 can achieve remarkable alignment with human judgment, particularly when assessing the thoroughness of responses.
 This paradigm shift suggests that AI systems themselves may offer the most effective tools for evaluating complex natural language generation tasks.
-LLMs have become widely adopted as evaluators in benchmarks like AlpacaEval [23], MT-Bench [13], and Chatbot Arena [18], where models such as GPT-4 conduct pairwise preference assessments based on helpfulness, factuality, and engagement.
-Frameworks like G-Eval [43] further enable fine-grained scoring by prompting models to assess specific dimensions using structured rubrics.
+LLMs have become widely adopted as evaluators in benchmarks like AlpacaEval [dubois2024length], MT-Bench [bai2024mt], and Chatbot Arena [chiang2024chatbot], where models such as GPT-4 conduct pairwise preference assessments based on helpfulness, factuality, and engagement.
+Frameworks like G-Eval [liu2023g] further enable fine-grained scoring by prompting models to assess specific dimensions using structured rubrics.
 These LLM-based evaluations have shown stronger alignment with human judgments than traditional metrics, especially on long-form and knowledge-intensive tasks.
@@
 D.1.2 Leveraging an Interpretable Ensemble of Reasoning LLMs
-To address the limitations of single-model, single-pass evaluation pipelines, we propose an interpretable and robust evaluation framework based on an ensemble of reasoning-capable LLMs: Deepseek-R1-Distilled [34], Qwen-3-32B [5], and Phi-4-Reasoning [6].
+To address the limitations of single-model, single-pass evaluation pipelines, we propose an interpretable and robust evaluation framework based on an ensemble of reasoning-capable LLMs: Deepseek-R1-Distilled [guo2025deepseek], Qwen-3-32B [githubQwen3Qwen3_Technical_ReportpdfMain], and Phi-4-Reasoning [abdin2025phi].
 These models were selected not only for their demonstrated strength in multi-hop reasoning and long-form comprehension but also for their open accessibility, ensuring that our evaluation protocol is fully transparent and reproducible without dependence on proprietary APIs.
@@
 Fleiss’ Kappa: To measure categorical agreement across multiple judges, we binarize evaluation outcomes (e.g., correct vs.
-incorrect) and compute Fleiss’ [28], a widely used metric for evaluating agreement on nominal data among fixed raters.
+incorrect) and compute Fleiss’ [fleiss1971measuring], a widely used metric for evaluating agreement on nominal data among fixed raters.
 This quantifies how consistently the LLM judges classify outputs beyond what would be expected by chance.
-Kendall’s W (Coefficient of Concordance: For tasks involving ordinal scoring or ranking of generations, we compute Kendall’s W [27], a non-parametric measure of rank correlation.
+Kendall’s W (Coefficient of Concordance: For tasks involving ordinal scoring or ranking of generations, we compute Kendall’s W [field2005k], a non-parametric measure of rank correlation.
 This accounts for judges using different scoring scales by focusing on relative orderings.
@@
 ICC is particularly suited for this setting as it accounts for both within-subject and between-subject variability, providing a continuous-scale assessment of intra-rater consistency.
-Following established interpretative guidelines [20] we classify ICC values into bands (e.g., moderate, good, excellent) to report the strength of reliability for each judge.
+Following established interpretative guidelines [cicchetti1994guidelines] we classify ICC values into bands (e.g., moderate, good, excellent) to report the strength of reliability for each judge.
 To assess the consistency of our LLM-based judges, we conducted a three-run intra-rater reliability analysis using the Intraclass Correlation Coefficient (ICC2) shown in Figure 7, separately computed for binary ID judgments and ordinal reasoning assessments.
@@
 We also observe that LLaVA-based models tend to show more intra-model fluctuation, potentially due to less structured or variable outputs.
-D.3 Diagnostic Parsimony
+D.3 Management (MG) Evaluation Criteria and W-Sum Score
+To evaluate model performance on the Management (MG) subset, we adopt a four-dimensional evaluation framework capturing complementary aspects of answer quality: Accuracy, Relevance, Completeness, and Parsimony.
+These dimensions were chosen based on findings from the long-form QA literature and insights from agricultural experts.
+Accuracy measures the factual correctness of the recommendation or diagnosis.
+Relevance assesses whether the response remains on-topic and directly addresses the user’s query.
+Completeness evaluates whether essential details (e.g., causal factors, conditions, or next-step actions) are covered.
+Parsimony rewards concise, evidence-based explanations that avoid unnecessary speculation or verbosity.
+The first three dimensions align with multi-axis evaluation principles in long-form QA studies [xu2023critical, beyondcortes2024beyond], which demonstrate that completeness and relevance correlate more strongly with human judgments than surface-level metrics such as ROUGE or BERTScore.
+The fourth dimension, Parsimony, is inspired by the cognitive and philosophical principle of simplicity (Occam’s razor), emphasizing clear and minimal explanations in expert communication.
+We provide a detailed discussion and illustrative examples of diagnostic parsimony in Appendix D.4.
+To enable model comparison with a single scalar, we compute a Weighted-Sum (W-Sum) score that aggregates these four dimensions.
+The adopted weight ratio of 2:1:1:1 reflects expert consensus that factual Accuracy should carry the greatest influence, since factual errors can lead to misleading or harmful recommendations.
+While we also report per-dimension results for transparency, the W-Sum provides a concise, interpretable measure for aggregate ranking.
+D.4 Diagnostic Parsimony
 Diagnostic parsimony refers to the principled restraint in offering explanations, favoring the simplest account consistent with observed evidence while avoiding unnecessary speculation.
-[45] shows that humans have a robust bias toward simpler explanations, not just for cognitive ease but because simplicity aids understanding, memory, and decision-making.
+[pacer2017ockham] shows that humans have a robust bias toward simpler explanations, not just for cognitive ease but because simplicity aids understanding, memory, and decision-making.
 In agriculture, this principle is especially critical: farmers and gardeners often seek immediate, actionable advice under time-sensitive or resource-constrained conditions.
@@
 LLM/LVLM answers in such knowledge-intensive domains must therefore strive to be concise yet comprehensive, ensuring the user gets just enough information to act confidently.
-Studies show that human evaluators, particularly non-experts, tend to prefer shorter, to-the-point responses, even when some details are omitted [65].
+Studies show that human evaluators, particularly non-experts, tend to prefer shorter, to-the-point responses, even when some details are omitted [xu2023critical].
 However, many instruction-following LLMs/LVLMs often default to verbose outputs, including tangential or overly cautious elaborations.
@@
 This behavior not only hinders user trust and satisfaction but also confounds traditional QA metrics like exact match (EM) and F1, which are poorly suited to judge informativeness under parsimony constraints.
-[9].
+[adlakha2024evaluating].
 Appendix E Additional MMST Benchmark Results
@@
 Consistent with the ID results, the proprietary GPT-4.1 model dominates most of metrics across all judges, outscoring the best open-source competitor by roughly absolute points in Accuracy and by – in the Relevance and Completeness.
-| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||||||||
+| Model | DeepSeek‑R1‑Distill | Qwen3‑32B | Phi‑4‑reasoning | |||||||||
 | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | |
@@
 | InternVL3-78B | 2.97 | |||||||||||
-Models are color-coded by type: closed-source models in red/light red, open-source models in blue/light blue.
-Scores represent performance on four key metrics (Accuracy / Relevance / Completeness / Parsimony) on a 0-4 scale.
+Models are color‑coded by type: closed‑source models in red/light red, open‑source models in blue/light blue.
+Scores represent performance on four key metrics (Accuracy / Relevance / Completeness / Parsimony) on a 0‑4 scale.
 Bold purple values highlight the best performance on each metric within each benchmark.
@@
 Utility of metadata.
-Tables 9 and 10 contrast model performance when the question is presented with versus without the user’s location and timestamp.
-For the Identification benchmark (Table 9), adding metadata yields only modest absolute gains ( pp ID%) and virtually no change in reasoning quality ().
-The strongest closed-source model (GPT-4.1) and the largest open-source Qwen-72B both extract a small benefit (+1.6 and +0.6 pp, respectively), whereas lighter counterparts (e.g., GPT-4.1-mini, Gemma-3B) see negligible or even negative shifts, suggesting that metadata can introduce noise unless the model possesses sufficient capacity to interpret it.
+Tables 9 and 10 compare model performance when each question is presented with versus without the user’s location and timestamp.
+Metadata is incorporated by appending this spatiotemporal context in natural-language form to the image–text input during inference.
+For the Identification benchmark, gains are modest (,pp ID%) with negligible changes in reasoning quality ().
+Stronger models (GPT-4.1, Qwen-72B) show slight improvements (+1.6 and +0.6,pp), whereas smaller ones (GPT-4.1-mini, Gemma-3B) show minimal or negative shifts, indicating that metadata can add noise without sufficient model capacity.
 A similar pattern emerges on the Management benchmark (Table 10): absolute deltas across accuracy, relevance, completeness, and parsimony remain within , yet the direction of change is revealing.
@@
 Hardware resources: one NVIDIA H200 GPU suffices for the QwenVL-2.5-3 B and 7 B models, while the 32 B model is trained on two H200 cards—enabling the full eight-epoch run.
-Effect of LoRA fine-tuning.
+Effect of LoRA fine‐tuning.
 Figure 10 tracks identification accuracy and reasoning accuracy on seen vs.
@@
 Identification requires the model to emit an explicit entity name; if that name never appeared in the fine-tuning set, fine-tuning does not help.
-E.5 Category-Wise Breakdown Results
-[…]
-E.5.1 Plant Identification Reults (MMST Standard)
-| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||
+Appendix F Error Analysis
+We conducted a quantitative error analysis to complement our main evaluation results for deeper insights into model failure modes.
+We performed a stratified manual error review across three scenarios: (i) GPT-4.1 failures ( of 2,395), (ii) Qwen2.5-VL-32B failures where GPT-4.1 succeeds ( of 1,078), and (iii) joint failures ( of 300).
+We focused on instances with zero identification accuracy and compared the failure modes and root causes between the proprietary GPT-4.1 and the open-source Qwen2.5-VL-32B.
+The observed errors cluster into two tiers: Fundamental Domain Challenges and Open-Source Model Systematic Failures.
+F.1 Tier 1: Fundamental Domain Challenges
+These represent the current boundaries of what even the most advanced LVLMs can achieve in agricultural reasoning and consultation contexts.
+Failures in this tier occur in GPT-4.1, the best-performing model in our evaluation, indicating inherent limitations of current LVLMs in agricultural expert reasoning.
+Percentages reflect the relative frequency of each error type within sampled failures
+Specialist Knowledge Gaps (35% of cases): These involve missing domain-specific understanding of biological phenomena requiring deep expertise beyond general training data.
+The LVLMs fail to recognize or reason about biological processes.
+Example: GPT-4.1 could not interpret “appleleaf blister mite” damage or its developmental process.
+Complex Diagnostic Reasoning (10% of cases): These involve correct symptom observation but incorrect causal inference or disease conclusion.
+Despite accurate visual analysis, the models fail to reason over underlying biological causality.
+Example: Observing insect damage but misdiagnosing the pest species.
+Finegrained visual understanding (55% of cases): Majority of the errors occured due to model’s inability to pick up finegrained visual cues and distinguishing factors among different adjacent biological species or part occlusions of the main entity in images.
+F.2 Tier 2: Open-Source Model Systematic Failures
+These failure types occur primarily in Qwen2.5-VL-32B but not in GPT-4.1, revealing systematic capability gaps between open-source and proprietary LVLMs.
+Vision-Language Integration Issues (40% of cases): Failures in integrating visual cues with domain-specific terminology lead to misidentification among visually similar species.
+Domain-specific finetuning mitigates this issue.
+Example: Qwen2.5-VL-32B confusing a longhorn beetle with a June beetle.
+Diagnostic Hedging Bias (25% of cases): The open-source model frequently exhibits hedging behavior, avoiding decisive expert-level conclusions.
+Generalist Reasoning Bias (20% of cases): The model applies broad, non-specific reasoning, listing multiple generic possibilities instead of providing a targeted diagnosis.
+F.3 Quantitative Error Pattern Summary
+| Failure Category | Percentage | Model |
+| Tier 1: Fundamental Domain Challenges (GPT-4.1 Failures) | ||
+| Specialist Knowledge Gaps | 35% | GPT-4.1 |
+| Complex Diagnostic Reasoning | 10% | GPT-4.1 |
+| Finegrained visual understanding | 55% | GPT-4.1 |
+| Tier 2: Open-Source Model Systematic Failures (Qwen2.5-VL-32B vs. GPT-4.1) | ||
+| Vision-Language Integration Issues | 40% | Qwen2.5-VL-32B |
+| Diagnostic Hedging Bias | 25% | Qwen2.5-VL-32B |
+| Generalist Reasoning Bias | 20% | Qwen2.5-VL-32B |
+Critical Finding: In joint failures, GPT-4.1 still demonstrates superior reasoning quality in 60% of cases, reflecting deeper integration of visual and textual cues and more expert-like reasoning patterns, even when incorrect in final identification.
+F.4 Performance on Common and Rare Entities
+The MIRAGE dataset exhibits a long-tail entity distribution, where models consistently perform worse on rare entities than on common ones.
+This imbalance exposes a central challenge for agricultural AI: models must recognize thousands of rare species that occur infrequently in training data.
+Addressing this limitation will require domain adaptation and long-tail learning approaches beyond standard large-scale training.
+| Model | DeepSeek-R1 | Qwen3-32B | Phi-4-reasoning |
+| GPT-4.1 | 53.1, 3.15 | 52.7, 2.84 | 52.2, 3.20 |
+| GPT-4.1-mini | 42.1, 2.82 | 41.3, 2.53 | 40.5, 2.95 |
+| Qwen2.5-VL-32B-Instruct | 32.9, 2.62 | 32.2, 2.23 | 31.6, 2.62 |
+| Qwen2.5-VL-72B-Instruct | 38.1, 2.65 | 38.4, 2.27 | 37.1, 2.63 |
+| Model | DeepSeek-R1 | Qwen3-32B | Phi-4-reasoning |
+| GPT-4.1 | 38.5, 3.01 | 38.9, 2.73 | 35.4, 3.14 |
+| GPT-4.1-mini | 32.1, 2.76 | 30.4, 2.50 | 26.8, 2.94 |
+| Qwen2.5-VL-32B-Instruct | 21.1, 2.47 | 20.4, 2.14 | 18.2, 2.53 |
+| Qwen2.5-VL-72B-Instruct | 25.5, 2.56 | 24.5, 2.18 | 22.1, 2.57 |
+Building on these findings, we outline a roadmap for advancing multimodal large vision-language models (VLMs) toward robust, interactive agricultural consultation systems.
+Future versions MIRAGE should integrate agentic capabilities that enable models to proactively utilize time- and location-based context—for example, adapting responses based on seasonal patterns, regional crop profiles, or current environmental conditions.
+This contextual grounding will allow evaluations to move beyond static reasoning and toward dynamic, situation-aware dialogue.
+To achieve this, future VLM development should emphasize the following directions:
+Contextual and Temporal Grounding: Incorporate temporal signals (e.g., crop cycles, weather timelines) and geospatial information (e.g., soil types, local pest occurrences) to support temporally and regionally coherent reasoning in consultation tasks.
+Interactive Conversational Reasoning: Extend MIRAGE’s multi-turn benchmark to simulate fully interactive dialogues, where models must perform clarification, conversational repair, and feedback incorporation, key aspects of real-world agricultural advisory systems.
+Domain-Specific Knowledge Integration: Bridge VLMs with structured agronomic knowledge graphs and expert-curated datasets to reduce specialist knowledge gaps identified in our error analysis.
+Adaptive Visual-Linguistic Understanding: Improve VLMs’ capacity to detect fine-grained visual cues and align them with domain terminology, addressing the vision-language integration failures seen in open-source models.
+Long-Tail Entity Recognition: Develop targeted fine-tuning and retrieval-augmented methods for handling rare species and diseases that dominate real-world agricultural problem distributions.
+These extensions will move MIRAGE and future VLMs from static perception and reasoning benchmarks toward fully interactive, contextually grounded expert systems.
+Such systems could ultimately assist human agricultural specialists in diagnosing issues, providing adaptive recommendations, and integrating multimodal data streams for decision support in diverse agricultural environments.
+F.5 Category-Wise Breakdown Results
+F.5.1 Plant Identification Reults (MMST Standard)
+| Model | DeepSeek‑R1‑Distill | Qwen3‑32B | Phi‑4‑reasoning | |||
 | Acc (%) | Reasoning | Acc (%) | Reasoning | Acc (%) | Reasoning | |
-| gpt-4.1 | 48.7 | 3.13 | 49.0 | 2.89 | 45.9 | 3.28 |
-| gpt-4.1-mini | ||||||
-| gpt-4o | ||||||
-| gpt-4o-mini | ||||||
-| claude-3-7-sonnet | ||||||
-| claude-3-5-sonnet | ||||||
-| claude-3-haiku | ||||||
-| Llama-4-Scout-17B-16E-Instruct | ||||||
-| llava-v1.6-mistral-7b-hf | ||||||
-| llava-onevision-qwen2-7b-ov-hf | ||||||
-| Qwen2.5-VL-3B-Instruct | ||||||
-| Qwen2.5-VL-7B-Instruct | ||||||
-| Qwen2.5-VL-32B-Instruct | ||||||
-| Qwen2.5-VL-72B-Instruct | ||||||
-| gemma-3-4b-it | ||||||
-| gemma-3-12b-it | ||||||
-| gemma-3-27b-it | ||||||
-| InternVL3-2B | ||||||
-| InternVL3-8B | ||||||
-| InternVL3-14B | ||||||
-| InternVL3-38B | ||||||
-| InternVL3-78B | ||||||
-Models are color-coded by type: closed-source models in red/light red, open-source models in blue/light blue. Bold purple values denote the best performance in each column.
-E.5.2 Insect and Pest Identification Results (MMST Standard)
-| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||
+| gpt‑4.1 | 48.7 | 3.13 | 49.0 | 2.89 | 45.9 | 3.28 |
+| gpt‑4.1‑mini | ||||||
+| gpt‑4o | ||||||
+| gpt‑4o‑mini | ||||||
+| claude‑3‑7‑sonnet | ||||||
+| claude‑3‑5‑sonnet | ||||||
+| claude‑3‑haiku | ||||||
+| Llama‑4‑Scout‑17B‑16E‑Instruct | ||||||
+| llava‑v1.6‑mistral‑7b‑hf | ||||||
+| llava‑onevision‑qwen2‑7b‑ov‑hf | ||||||
+| Qwen2.5‑VL‑3B‑Instruct | ||||||
+| Qwen2.5‑VL‑7B‑Instruct | ||||||
+| Qwen2.5‑VL‑32B‑Instruct | ||||||
+| Qwen2.5‑VL‑72B‑Instruct | ||||||
+| gemma‑3‑4b‑it | ||||||
+| gemma‑3‑12b‑it | ||||||
+| gemma‑3‑27b‑it | ||||||
+| InternVL3‑2B | ||||||
+| InternVL3‑8B | ||||||
+| InternVL3‑14B | ||||||
+| InternVL3‑38B | ||||||
+| InternVL3‑78B | ||||||
+Models are color‑coded by type: closed‑source models in red/light red, open‑source models in blue/light blue. Bold purple values denote the best performance in each column.
+F.5.2 Insect and Pest Identification Results (MMST Standard)
+| Model | DeepSeek‑R1‑Distill | Qwen3‑32B | Phi‑4‑reasoning | |||
 | Acc (%) | Reasoning | Acc (%) | Reasoning | Acc (%) | Reasoning | |
-| gpt-4.1 | 34.9 | 2.93 | 34.9 | 2.60 | 32.6 | 3.01 |
-| gpt-4.1-mini | ||||||
-| gpt-4o | ||||||
-| gpt-4o-mini | ||||||
-| claude-3-7-sonnet | ||||||
-| claude-3-5-sonnet | ||||||
-| claude-3-haiku | ||||||
-| Llama-4-Scout-17B-16E-Instruct | ||||||
-| llava-v1.6-mistral-7b-hf | ||||||
-| llava-onevision-qwen2-7b-ov-hf | ||||||
-| Qwen2.5-VL-3B-Instruct | ||||||
-| Qwen2.5-VL-7B-Instruct | ||||||
-| Qwen2.5-VL-32B-Instruct | ||||||
-| Qwen2.5-VL-72B-Instruct | ||||||
-| gemma-3-4b-it | ||||||
-| gemma-3-12b-it | ||||||
-| gemma-3-27b-it | ||||||
-| InternVL3-2B | ||||||
-| InternVL3-8B | ||||||
-| InternVL3-14B | ||||||
-| InternVL3-38B | ||||||
-| InternVL3-78B | ||||||
-Models are color-coded by type: closed-source models in red/light red, open-source models in blue/light blue. Bold purple values denote the best performance in each column.
-E.5.3 Plant Disease Identification Results (MMST Standard)
-| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||
+| gpt‑4.1 | 34.9 | 2.93 | 34.9 | 2.60 | 32.6 | 3.01 |
+| gpt‑4.1‑mini | ||||||
+| gpt‑4o | ||||||
+| gpt‑4o‑mini | ||||||
+| claude‑3‑7‑sonnet | ||||||
+| claude‑3‑5‑sonnet | ||||||
+| claude‑3‑haiku | ||||||
+| Llama‑4‑Scout‑17B‑16E‑Instruct | ||||||
+| llava‑v1.6‑mistral‑7b‑hf | ||||||
+| llava‑onevision‑qwen2‑7b‑ov‑hf | ||||||
+| Qwen2.5‑VL‑3B‑Instruct | ||||||
+| Qwen2.5‑VL‑7B‑Instruct | ||||||
+| Qwen2.5‑VL‑32B‑Instruct | ||||||
+| Qwen2.5‑VL‑72B‑Instruct | ||||||
+| gemma‑3‑4b‑it | ||||||
+| gemma‑3‑12b‑it | ||||||
+| gemma‑3‑27b‑it | ||||||
+| InternVL3‑2B | ||||||
+| InternVL3‑8B | ||||||
+| InternVL3‑14B | ||||||
+| InternVL3‑38B | ||||||
+| InternVL3‑78B | ||||||
+Models are color‑coded by type: closed‑source models in red/light red, open‑source models in blue/light blue. Bold purple values denote the best performance in each column.
+F.5.3 Plant Disease Identification Results (MMST Standard)
+| Model | DeepSeek‑R1‑Distill | Qwen3‑32B | Phi‑4‑reasoning | |||
 | Acc (%) | Reasoning | Acc (%) | Reasoning | Acc (%) | Reasoning | |
-| gpt-4.1 | 45.7 | 3.06 | 44.6 | 2.62 | 46.0 | 2.97 |
-| gpt-4.1-mini | ||||||
-| gpt-4o | ||||||
-| gpt-4o-mini | ||||||
-| claude-3-7-sonnet | ||||||
-| claude-3-5-sonnet | ||||||
-| claude-3-haiku | ||||||
-| Llama-4-Scout-17B-16E-Instruct | ||||||
-| llava-v1.6-mistral-7b-hf | ||||||
-| llava-onevision-qwen2-7b-ov-hf | ||||||
-| Qwen2.5-VL-3B-Instruct | ||||||
-| Qwen2.5-VL-7B-Instruct | ||||||
-| Qwen2.5-VL-32B-Instruct | ||||||
-| Qwen2.5-VL-72B-Instruct | ||||||
-| gemma-3-4b-it | ||||||
-| gemma-3-12b-it | ||||||
-| gemma-3-27b-it | ||||||
-| InternVL3-2B | ||||||
-| InternVL3-8B | ||||||
-| InternVL3-14B | ||||||
-| InternVL3-38B | ||||||
-| InternVL3-78B | ||||||
-Models are color-coded by type: closed-source models in red/light red, open-source models in blue/light blue. Bold purple values denote the best performance in each column.
-E.5.4 Plant Disease Management Results (MMST Standard)
-| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||||||||
+| gpt‑4.1 | 45.7 | 3.06 | 44.6 | 2.62 | 46.0 | 2.97 |
+| gpt‑4.1‑mini | ||||||
+| gpt‑4o | ||||||
+| gpt‑4o‑mini | ||||||
+| claude‑3‑7‑sonnet | ||||||
+| claude‑3‑5‑sonnet | ||||||
+| claude‑3‑haiku | ||||||
+| Llama‑4‑Scout‑17B‑16E‑Instruct | ||||||
+| llava‑v1.6‑mistral‑7b‑hf | ||||||
+| llava‑onevision‑qwen2‑7b‑ov‑hf | ||||||
+| Qwen2.5‑VL‑3B‑Instruct | ||||||
+| Qwen2.5‑VL‑7B‑Instruct | ||||||
+| Qwen2.5‑VL‑32B‑Instruct | ||||||
+| Qwen2.5‑VL‑72B‑Instruct | ||||||
+| gemma‑3‑4b‑it | ||||||
+| gemma‑3‑12b‑it | ||||||
+| gemma‑3‑27b‑it | ||||||
+| InternVL3‑2B | ||||||
+| InternVL3‑8B | ||||||
+| InternVL3‑14B | ||||||
+| InternVL3‑38B | ||||||
+| InternVL3‑78B | ||||||
+Models are color‑coded by type: closed‑source models in red/light red, open‑source models in blue/light blue. Bold purple values denote the best performance in each column.
+F.5.4 Plant Disease Management Results (MMST Standard)
+| Model | DeepSeek‑R1‑Distill | Qwen3‑32B | Phi‑4‑reasoning | |||||||||
 | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | |
-| gpt-4.1 | 3.05 | 3.54 | 3.31 | 2.87 | 3.17 | 3.54 | 2.90 | 2.92 | 3.15 | 3.53 | 3.12 | 3.05 |
-| gpt-4.1-mini | ||||||||||||
-| gpt-4o | ||||||||||||
-| gpt-4o-mini | ||||||||||||
-| claude-3-7-sonnet | ||||||||||||
-| claude-3-5-sonnet | ||||||||||||
-| claude-3-haiku | ||||||||||||
-| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
-| llava-v1.6-mistral-7b-hf | ||||||||||||
-| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
-| Qwen2.5-VL-3B-Instruct | ||||||||||||
-| Qwen2.5-VL-7B-Instruct | ||||||||||||
-| Qwen2.5-VL-32B-Instruct | ||||||||||||
-| Qwen2.5-VL-72B-Instruct | ||||||||||||
-| gemma-3-4b-it | ||||||||||||
-| gemma-3-12b-it | ||||||||||||
-| gemma-3-27b-it | ||||||||||||
-| InternVL3-2B | ||||||||||||
-| InternVL3-8B | ||||||||||||
-| InternVL3-14B | ||||||||||||
-| InternVL3-38B | ||||||||||||
-| InternVL3-78B | ||||||||||||
-Models are color-coded by type: closed-source models in red/light red, open-source models in blue/light blue.
+| gpt‑4.1 | 3.05 | 3.54 | 3.31 | 2.87 | 3.17 | 3.54 | 2.90 | 2.92 | 3.15 | 3.53 | 3.12 | 3.05 |
+| gpt‑4.1‑mini | ||||||||||||
+| gpt‑4o | ||||||||||||
+| gpt‑4o‑mini | ||||||||||||
+| claude‑3‑7‑sonnet | ||||||||||||
+| claude‑3‑5‑sonnet | ||||||||||||
+| claude‑3‑haiku | ||||||||||||
+| Llama‑4‑Scout‑17B‑16E‑Instruct | ||||||||||||
+| llava‑v1.6‑mistral‑7b‑hf | ||||||||||||
+| llava‑onevision‑qwen2‑7b‑ov‑hf | ||||||||||||
+| Qwen2.5‑VL‑3B‑Instruct | ||||||||||||
+| Qwen2.5‑VL‑7B‑Instruct | ||||||||||||
+| Qwen2.5‑VL‑32B‑Instruct | ||||||||||||
+| Qwen2.5‑VL‑72B‑Instruct | ||||||||||||
+| gemma‑3‑4b‑it | ||||||||||||
+| gemma‑3‑12b‑it | ||||||||||||
+| gemma‑3‑27b‑it | ||||||||||||
+| InternVL3‑2B | ||||||||||||
+| InternVL3‑8B | ||||||||||||
+| InternVL3‑14B | ||||||||||||
+| InternVL3‑38B | ||||||||||||
+| InternVL3‑78B | ||||||||||||
+Models are color‑coded by type: closed‑source models in red/light red, open‑source models in blue/light blue.
 Scores are given on a 0–4 scale for Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars).
 Bold purple numbers denote the best performance for each metric within a column block.
-E.5.5 Insect and Pest Management Results (MMST Standard)
-| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||||||||
+F.5.5 Insect and Pest Management Results (MMST Standard)
+| Model | DeepSeek‑R1‑Distill | Qwen3‑32B | Phi‑4‑reasoning | |||||||||
 | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | |
-| gpt-4.1 | 3.05 | 3.51 | 3.33 | 2.87 | 3.18 | 3.53 | 2.98 | 2.98 | 3.11 | 3.52 | 3.15 | 3.06 |
-| gpt-4.1-mini | ||||||||||||
-| gpt-4o | ||||||||||||
-| gpt-4o-mini | ||||||||||||
-| nova-pro | ||||||||||||
-| claude-3-7-sonnet | ||||||||||||
-| claude-3-5-sonnet | ||||||||||||
-| claude-3-haiku | ||||||||||||
-| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
-| llava-v1.6-mistral-7b-hf | ||||||||||||
-| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
-| Qwen2.5-VL-3B-Instruct | ||||||||||||
-| Qwen2.5-VL-7B-Instruct | ||||||||||||
-| Qwen2.5-VL-32B-Instruct | ||||||||||||
-| Qwen2.5-VL-72B-Instruct | ||||||||||||
-| gemma-3-4b-it | ||||||||||||
-| gemma-3-12b-it | ||||||||||||
-| gemma-3-27b-it | ||||||||||||
-| InternVL3-2B | ||||||||||||
-| InternVL3-8B | ||||||||||||
-| InternVL3-14B | ||||||||||||
-| InternVL3-38B | ||||||||||||
-| InternVL3-78B | ||||||||||||
-Models are color-coded by type: closed-source (red) and open-source (blue).
-Scores range from 0-4 for Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars).
+| gpt‑4.1 | 3.05 | 3.51 | 3.33 | 2.87 | 3.18 | 3.53 | 2.98 | 2.98 | 3.11 | 3.52 | 3.15 | 3.06 |
+| gpt‑4.1‑mini | ||||||||||||
+| gpt‑4o | ||||||||||||
+| gpt‑4o‑mini | ||||||||||||
+| nova‑pro | ||||||||||||
+| claude‑3‑7‑sonnet | ||||||||||||
+| claude‑3‑5‑sonnet | ||||||||||||
+| claude‑3‑haiku | ||||||||||||
+| Llama‑4‑Scout‑17B‑16E‑Instruct | ||||||||||||
+| llava‑v1.6‑mistral‑7b‑hf | ||||||||||||
+| llava‑onevision‑qwen2‑7b‑ov‑hf | ||||||||||||
+| Qwen2.5‑VL‑3B‑Instruct | ||||||||||||
+| Qwen2.5‑VL‑7B‑Instruct | ||||||||||||
+| Qwen2.5‑VL‑32B‑Instruct | ||||||||||||
+| Qwen2.5‑VL‑72B‑Instruct | ||||||||||||
+| gemma‑3‑4b‑it | ||||||||||||
+| gemma‑3‑12b‑it | ||||||||||||
+| gemma‑3‑27b‑it | ||||||||||||
+| InternVL3‑2B | ||||||||||||
+| InternVL3‑8B | ||||||||||||
+| InternVL3‑14B | ||||||||||||
+| InternVL3‑38B | ||||||||||||
+| InternVL3‑78B | ||||||||||||
+Models are color‑coded by type: closed‑source (red) and open‑source (blue).
+Scores range from 0‑4 for Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars).
 Bold purple indicates the best score in each column block.
-E.5.6 Plant Care and Gardening Guidance Results (MMST Standard)
-| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||||||||
+F.5.6 Plant Care and Gardening Guidance Results (MMST Standard)
+| Model | DeepSeek‑R1‑Distill | Qwen3‑32B | Phi‑4‑reasoning | |||||||||
 | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | |
-| gpt-4.1 | 3.38 | 3.73 | 3.55 | 3.01 | 3.45 | 3.73 | 3.17 | 3.12 | 3.52 | 3.80 | 3.43 | 3.24 |
-| gpt-4.1-mini | ||||||||||||
-| gpt-4o | ||||||||||||
-| gpt-4o-mini | ||||||||||||
-| nova-pro | ||||||||||||
-| claude-3-7-sonnet | ||||||||||||
-| claude-3-5-sonnet | ||||||||||||
-| claude-3-haiku | ||||||||||||
-| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
-| llava-v1.6-mistral-7b-hf | ||||||||||||
-| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
-| Qwen2.5-VL-3B-Instruct | ||||||||||||
-| Qwen2.5-VL-7B-Instruct | ||||||||||||
-| Qwen2.5-VL-32B-Instruct | ||||||||||||
-| Qwen2.5-VL-72B-Instruct | ||||||||||||
-| gemma-3-4b-it | ||||||||||||
-| gemma-3-12b-it | ||||||||||||
-| gemma-3-27b-it | ||||||||||||
-| InternVL3-2B | ||||||||||||
-| InternVL3-8B | ||||||||||||
-| InternVL3-14B | ||||||||||||
-| InternVL3-38B | ||||||||||||
-| InternVL3-78B | ||||||||||||
-Models are color-coded by type: closed-source (red) and open-source (blue).
+| gpt‑4.1 | 3.38 | 3.73 | 3.55 | 3.01 | 3.45 | 3.73 | 3.17 | 3.12 | 3.52 | 3.80 | 3.43 | 3.24 |
+| gpt‑4.1‑mini | ||||||||||||
+| gpt‑4o | ||||||||||||
+| gpt‑4o‑mini | ||||||||||||
+| nova‑pro | ||||||||||||
+| claude‑3‑7‑sonnet | ||||||||||||
+| claude‑3‑5‑sonnet | ||||||||||||
+| claude‑3‑haiku | ||||||||||||
+| Llama‑4‑Scout‑17B‑16E‑Instruct | ||||||||||||
+| llava‑v1.6‑mistral‑7b‑hf | ||||||||||||
+| llava‑onevision‑qwen2‑7b‑ov‑hf | ||||||||||||
+| Qwen2.5‑VL‑3B‑Instruct | ||||||||||||
+| Qwen2.5‑VL‑7B‑Instruct | ||||||||||||
+| Qwen2.5‑VL‑32B‑Instruct | ||||||||||||
+| Qwen2.5‑VL‑72B‑Instruct | ||||||||||||
+| gemma‑3‑4b‑it | ||||||||||||
+| gemma‑3‑12b‑it | ||||||||||||
+| gemma‑3‑27b‑it | ||||||||||||
+| InternVL3‑2B | ||||||||||||
+| InternVL3‑8B | ||||||||||||
+| InternVL3‑14B | ||||||||||||
+| InternVL3‑38B | ||||||||||||
+| InternVL3‑78B | ||||||||||||
+Models are color‑coded by type: closed‑source (red) and open‑source (blue).
 Each metric is scored on a 0–4 scale: Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars).
 Bold purple values indicate the best performance within each column block.
-E.5.7 Weeds/Invasive Plants Management Results (MMST Standard)
-| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||||||||
+F.5.7 Weeds/Invasive Plants Management Results (MMST Standard)
+| Model | DeepSeek‑R1‑Distill | Qwen3‑32B | Phi‑4‑reasoning | |||||||||
 | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | |
-| gpt-4.1 | 2.90 | 3.41 | 3.14 | 3.01 | 3.37 | 2.87 | 2.89 | 3.40 | 3.00 | |||
-| gpt-4.1-mini | ||||||||||||
-| gpt-4o | 2.95 | 3.17 | 3.11 | |||||||||
-| gpt-4o-mini | ||||||||||||
-| nova-pro | ||||||||||||
-| claude-3-7-sonnet | ||||||||||||
-| claude-3-5-sonnet | ||||||||||||
-| claude-3-haiku | ||||||||||||
-| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
-| llava-v1.6-mistral-7b-hf | ||||||||||||
-| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
-| Qwen2.5-VL-3B-Instruct | ||||||||||||
-| Qwen2.5-VL-7B-Instruct | ||||||||||||
-| Qwen2.5-VL-32B-Instruct | ||||||||||||
-| Qwen2.5-VL-72B-Instruct | ||||||||||||
-| gemma-3-4b-it | ||||||||||||
-| gemma-3-12b-it | ||||||||||||
-| gemma-3-27b-it | ||||||||||||
-| InternVL3-2B | ||||||||||||
-| InternVL3-8B | ||||||||||||
-| InternVL3-14B | ||||||||||||
-| InternVL3-38B | ||||||||||||
-| InternVL3-78B | ||||||||||||
-Models are color-coded by type: closed-source (red) and open-source (blue).
+| gpt‑4.1 | 2.90 | 3.41 | 3.14 | 3.01 | 3.37 | 2.87 | 2.89 | 3.40 | 3.00 | |||
+| gpt‑4.1‑mini | ||||||||||||
+| gpt‑4o | 2.95 | 3.17 | 3.11 | |||||||||
+| gpt‑4o‑mini | ||||||||||||
+| nova‑pro | ||||||||||||
+| claude‑3‑7‑sonnet | ||||||||||||
+| claude‑3‑5‑sonnet | ||||||||||||
+| claude‑3‑haiku | ||||||||||||
+| Llama‑4‑Scout‑17B‑16E‑Instruct | ||||||||||||
+| llava‑v1.6‑mistral‑7b‑hf | ||||||||||||
+| llava‑onevision‑qwen2‑7b‑ov‑hf | ||||||||||||
+| Qwen2.5‑VL‑3B‑Instruct | ||||||||||||
+| Qwen2.5‑VL‑7B‑Instruct | ||||||||||||
+| Qwen2.5‑VL‑32B‑Instruct | ||||||||||||
+| Qwen2.5‑VL‑72B‑Instruct | ||||||||||||
+| gemma‑3‑4b‑it | ||||||||||||
+| gemma‑3‑12b‑it | ||||||||||||
+| gemma‑3‑27b‑it | ||||||||||||
+| InternVL3‑2B | ||||||||||||
+| InternVL3‑8B | ||||||||||||
+| InternVL3‑14B | ||||||||||||
+| InternVL3‑38B | ||||||||||||
+| InternVL3‑78B | ||||||||||||
+Models are color‑coded by type: closed‑source (red) and open‑source (blue).
 Scores range from 0–4 across Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars).
 Bold purple marks the highest score for each metric within a column block.
-E.5.8 Plant Disease Management Results (MMST Contextual)
-| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||||||||
+F.5.8 Plant Disease Management Results (MMST Contextual)
+| Model | DeepSeek‑R1‑Distill | Qwen3‑32B | Phi‑4‑reasoning | |||||||||
 | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | |
-| gpt-4.1 | 3.07 | 3.52 | 3.34 | 2.86 | 3.16 | 3.49 | 2.91 | 3.09 | 3.44 | 3.08 | 2.96 | |
-| gpt-4.1-mini | ||||||||||||
-| gpt-4o | 3.17 | |||||||||||
-| gpt-4o-mini | ||||||||||||
-| nova-pro | ||||||||||||
-| claude-3-7-sonnet | ||||||||||||
-| claude-3-5-sonnet | ||||||||||||
-| claude-3-haiku | ||||||||||||
-| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
-| llava-v1.6-mistral-7b-hf | ||||||||||||
-| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
-| Qwen2.5-VL-3B-Instruct | ||||||||||||
-| Qwen2.5-VL-7B-Instruct | ||||||||||||
-| Qwen2.5-VL-32B-Instruct | ||||||||||||
-| Qwen2.5-VL-72B-Instruct | ||||||||||||
-| gemma-3-4b-it | ||||||||||||
-| gemma-3-12b-it | ||||||||||||
-| gemma-3-27b-it | ||||||||||||
-| InternVL3-2B | ||||||||||||
-| InternVL3-8B | ||||||||||||
-| InternVL3-14B | ||||||||||||
-| InternVL3-38B | ||||||||||||
-| InternVL3-78B | ||||||||||||
-Models are color-coded by type: closed-source (red) and open-source (blue).
+| gpt‑4.1 | 3.07 | 3.52 | 3.34 | 2.86 | 3.16 | 3.49 | 2.91 | 3.09 | 3.44 | 3.08 | 2.96 | |
+| gpt‑4.1‑mini | ||||||||||||
+| gpt‑4o | 3.17 | |||||||||||
+| gpt‑4o‑mini | ||||||||||||
+| nova‑pro | ||||||||||||
+| claude‑3‑7‑sonnet | ||||||||||||
+| claude‑3‑5‑sonnet | ||||||||||||
+| claude‑3‑haiku | ||||||||||||
+| Llama‑4‑Scout‑17B‑16E‑Instruct | ||||||||||||
+| llava‑v1.6‑mistral‑7b‑hf | ||||||||||||
+| llava‑onevision‑qwen2‑7b‑ov‑hf | ||||||||||||
+| Qwen2.5‑VL‑3B‑Instruct | ||||||||||||
+| Qwen2.5‑VL‑7B‑Instruct | ||||||||||||
+| Qwen2.5‑VL‑32B‑Instruct | ||||||||||||
+| Qwen2.5‑VL‑72B‑Instruct | ||||||||||||
+| gemma‑3‑4b‑it | ||||||||||||
+| gemma‑3‑12b‑it | ||||||||||||
+| gemma‑3‑27b‑it | ||||||||||||
+| InternVL3‑2B | ||||||||||||
+| InternVL3‑8B | ||||||||||||
+| InternVL3‑14B | ||||||||||||
+| InternVL3‑38B | ||||||||||||
+| InternVL3‑78B | ||||||||||||
+Models are color‑coded by type: closed‑source (red) and open‑source (blue).
 Scores range from 0–4 across Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars).
 Bold purple marks the highest score for each metric within a column block.
-E.5.9 Insect and Pest Management Results (MMST Contextual)
-| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||||||||
+F.5.9 Insect and Pest Management Results (MMST Contextual)
+| Model | DeepSeek‑R1‑Distill | Qwen3‑32B | Phi‑4‑reasoning | |||||||||
 | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | |
-| gpt-4.1 | 3.14 | 3.59 | 3.32 | 2.99 | 3.25 | 3.61 | 2.99 | 3.06 | 3.22 | 3.60 | 3.20 | 3.21 |
-| gpt-4.1-mini | ||||||||||||
-| gpt-4o | ||||||||||||
-| gpt-4o-mini | ||||||||||||
-| nova-pro | ||||||||||||
-| claude-3-7-sonnet | ||||||||||||
-| claude-3-5-sonnet | ||||||||||||
-| claude-3-haiku | ||||||||||||
-| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
-| llava-v1.6-mistral-7b-hf | ||||||||||||
-| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
-| Qwen2.5-VL-3B-Instruct | ||||||||||||
-| Qwen2.5-VL-7B-Instruct | ||||||||||||
-| Qwen2.5-VL-32B-Instruct | ||||||||||||
-| Qwen2.5-VL-72B-Instruct | ||||||||||||
-| gemma-3-4b-it | ||||||||||||
-| gemma-3-12b-it | ||||||||||||
-| gemma-3-27b-it | ||||||||||||
-| InternVL3-2B | ||||||||||||
-| InternVL3-8B | ||||||||||||
-| InternVL3-14B | ||||||||||||
-| InternVL3-38B | ||||||||||||
-| InternVL3-78B | ||||||||||||
-Models are color-coded by type: closed-source (red) and open-source (blue).
+| gpt‑4.1 | 3.14 | 3.59 | 3.32 | 2.99 | 3.25 | 3.61 | 2.99 | 3.06 | 3.22 | 3.60 | 3.20 | 3.21 |
+| gpt‑4.1‑mini | ||||||||||||
+| gpt‑4o | ||||||||||||
+| gpt‑4o‑mini | ||||||||||||
+| nova‑pro | ||||||||||||
+| claude‑3‑7‑sonnet | ||||||||||||
+| claude‑3‑5‑sonnet | ||||||||||||
+| claude‑3‑haiku | ||||||||||||
+| Llama‑4‑Scout‑17B‑16E‑Instruct | ||||||||||||
+| llava‑v1.6‑mistral‑7b‑hf | ||||||||||||
+| llava‑onevision‑qwen2‑7b‑ov‑hf | ||||||||||||
+| Qwen2.5‑VL‑3B‑Instruct | ||||||||||||
+| Qwen2.5‑VL‑7B‑Instruct | ||||||||||||
+| Qwen2.5‑VL‑32B‑Instruct | ||||||||||||
+| Qwen2.5‑VL‑72B‑Instruct | ||||||||||||
+| gemma‑3‑4b‑it | ||||||||||||
+| gemma‑3‑12b‑it | ||||||||||||
+| gemma‑3‑27b‑it | ||||||||||||
+| InternVL3‑2B | ||||||||||||
+| InternVL3‑8B | ||||||||||||
+| InternVL3‑14B | ||||||||||||
+| InternVL3‑38B | ||||||||||||
+| InternVL3‑78B | ||||||||||||
+Models are color‑coded by type: closed‑source (red) and open‑source (blue).
 Scores range from 0–4 for Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars).
 Bold purple highlights the best score for each metric within a column block.
-E.5.10 Plant Care and Gardening Guidance Results (MMST Contextual)
-| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||||||||
+F.5.10 Plant Care and Gardening Guidance Results (MMST Contextual)
+| Model | DeepSeek‑R1‑Distill | Qwen3‑32B | Phi‑4‑reasoning | |||||||||
 | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | |
-| gpt-4.1 | 3.40 | 3.73 | 3.54 | 3.46 | 3.72 | 3.16 | 3.53 | 3.80 | 3.42 | 3.27 | ||
-| gpt-4.1-mini | ||||||||||||
-| gpt-4o | 3.08 | |||||||||||
-| gpt-4o-mini | ||||||||||||
-| nova-pro | ||||||||||||
-| claude-3-7-sonnet | ||||||||||||
-| claude-3-5-sonnet | ||||||||||||
-| claude-3-haiku | ||||||||||||
-| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
-| llava-v1.6-mistral-7b-hf | ||||||||||||
-| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
-| Qwen2.5-VL-3B-Instruct | ||||||||||||
-| Qwen2.5-VL-7B-Instruct | ||||||||||||
-| Qwen2.5-VL-32B-Instruct | ||||||||||||
-| Qwen2.5-VL-72B-Instruct | ||||||||||||
-| gemma-3-4b-it | ||||||||||||
-| gemma-3-12b-it | ||||||||||||
-| gemma-3-27b-it | ||||||||||||
-| InternVL3-2B | ||||||||||||
-| InternVL3-8B | ||||||||||||
-| InternVL3-14B | ||||||||||||
-| InternVL3-38B | ||||||||||||
-| InternVL3-78B | 3.50 | |||||||||||
-Models are color-coded by type: closed-source (red) and open-source (blue).
+| gpt‑4.1 | 3.40 | 3.73 | 3.54 | 3.46 | 3.72 | 3.16 | 3.53 | 3.80 | 3.42 | 3.27 | ||
+| gpt‑4.1‑mini | ||||||||||||
+| gpt‑4o | 3.08 | |||||||||||
+| gpt‑4o‑mini | ||||||||||||
+| nova‑pro | ||||||||||||
+| claude‑3‑7‑sonnet | ||||||||||||
+| claude‑3‑5‑sonnet | ||||||||||||
+| claude‑3‑haiku | ||||||||||||
+| Llama‑4‑Scout‑17B‑16E‑Instruct | ||||||||||||
+| llava‑v1.6‑mistral‑7b‑hf | ||||||||||||
+| llava‑onevision‑qwen2‑7b‑ov‑hf | ||||||||||||
+| Qwen2.5‑VL‑3B‑Instruct | ||||||||||||
+| Qwen2.5‑VL‑7B‑Instruct | ||||||||||||
+| Qwen2.5‑VL‑32B‑Instruct | ||||||||||||
+| Qwen2.5‑VL‑72B‑Instruct | ||||||||||||
+| gemma‑3‑4b‑it | ||||||||||||
+| gemma‑3‑12b‑it | ||||||||||||
+| gemma‑3‑27b‑it | ||||||||||||
+| InternVL3‑2B | ||||||||||||
+| InternVL3‑8B | ||||||||||||
+| InternVL3‑14B | ||||||||||||
+| InternVL3‑38B | ||||||||||||
+| InternVL3‑78B | 3.50 | |||||||||||
+Models are color‑coded by type: closed‑source (red) and open‑source (blue).
 Scores range from 0–4 for Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars).
 Bold purple highlights the best score for each metric within a column block.
-E.5.11 Weeds/Invasive Plants Management Results (MMST Contextual)
-| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||||||||
+F.5.11 Weeds/Invasive Plants Management Results (MMST Contextual)
+| Model | DeepSeek‑R1‑Distill | Qwen3‑32B | Phi‑4‑reasoning | |||||||||
 | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | Acc | Rel | Comp | Pars | |
-| gpt-4.1 | 3.03 | 3.53 | 3.27 | 2.92 | 3.12 | 3.46 | 2.92 | 3.03 | 3.02 | 3.45 | 3.07 | 3.20 |
-| gpt-4.1-mini | ||||||||||||
-| gpt-4o | ||||||||||||
-| gpt-4o-mini | ||||||||||||
-| nova-pro | ||||||||||||
-| claude-3-7-sonnet | ||||||||||||
-| claude-3-5-sonnet | ||||||||||||
-| claude-3-haiku | ||||||||||||
-| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
-| llava-v1.6-mistral-7b-hf | ||||||||||||
-| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
-| Qwen2.5-VL-3B-Instruct | ||||||||||||
-| Qwen2.5-VL-7B-Instruct | ||||||||||||
-| Qwen2.5-VL-32B-Instruct | ||||||||||||
-| Qwen2.5-VL-72B-Instruct | ||||||||||||
-| gemma-3-4b-it | ||||||||||||
-| gemma-3-12b-it | ||||||||||||
-| gemma-3-27b-it | ||||||||||||
-| InternVL3-2B | ||||||||||||
-| InternVL3-8B | ||||||||||||
-| InternVL3-14B | ||||||||||||
-| InternVL3-38B | ||||||||||||
-| InternVL3-78B | ||||||||||||
-Models are color-coded by type: closed-source (red) and open-source (blue).
+| gpt‑4.1 | 3.03 | 3.53 | 3.27 | 2.92 | 3.12 | 3.46 | 2.92 | 3.03 | 3.02 | 3.45 | 3.07 | 3.20 |
+| gpt‑4.1‑mini | ||||||||||||
+| gpt‑4o | ||||||||||||
+| gpt‑4o‑mini | ||||||||||||
+| nova‑pro | ||||||||||||
+| claude‑3‑7‑sonnet | ||||||||||||
+| claude‑3‑5‑sonnet | ||||||||||||
+| claude‑3‑haiku | ||||||||||||
+| Llama‑4‑Scout‑17B‑16E‑Instruct | ||||||||||||
+| llava‑v1.6‑mistral‑7b‑hf | ||||||||||||
+| llava‑onevision‑qwen2‑7b‑ov‑hf | ||||||||||||
+| Qwen2.5‑VL‑3B‑Instruct | ||||||||||||
+| Qwen2.5‑VL‑7B‑Instruct | ||||||||||||
+| Qwen2.5‑VL‑32B‑Instruct | ||||||||||||
+| Qwen2.5‑VL‑72B‑Instruct | ||||||||||||
+| gemma‑3‑4b‑it | ||||||||||||
+| gemma‑3‑12b‑it | ||||||||||||
+| gemma‑3‑27b‑it | ||||||||||||
+| InternVL3‑2B | ||||||||||||
+| InternVL3‑8B | ||||||||||||
+| InternVL3‑14B | ||||||||||||
+| InternVL3‑38B | ||||||||||||
+| InternVL3‑78B | ||||||||||||
+Models are color‑coded by type: closed‑source (red) and open‑source (blue).
 Scores range from 0–4 for Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars).
 Bold purple marks the highest score for each metric within a column block.
-Appendix F MIRAGE-MMMT
-F.1 Benchmark Details
+Appendix G MIRAGE-MMMT
+G.1 Benchmark Details
 The MIRAGE-MMMT dataset as shown in Table 25, contains 861 multi-turn samples, each annotated with a high-level decision label—either Clarify (56.6%) or Respond (43.4%)—reflecting the expert’s intent in continuing the consultation.
@@
 These characteristics make the dataset well-suited for studying decision-making, goal inference, and clarification strategies in visually grounded, expert-guided dialogues.
-F.2 Task Definition
+G.2 Task Definition
 MIRAGE-MT is a multimodal, multi-turn benchmark designed to evaluate conversational expert agents in a consultative decision-making setting.
 Given a multi-turn dialogue and associated image(s) as shown in Figure 11, the agent must decide whether to ask a clarification question or provide a helpful response, and then generate the corresponding utterance.
-F.2.1 Input
+G.2.1 Input
 Let a multi-turn dialogue context be represented as a sequence:
@@
 which may provide visual context necessary for interpretation (e.g., pest damage, plant structure).
-F.2.2 Output
+G.2.2 Output
 The model must jointly predict:
@@
 If , then is an expert answer
-F.2.3 Goal Inference and Decision Policy
+G.2.3 Goal Inference and Decision Policy
 Let denote the user’s underlying goal (e.g., identifying a plant disease, choosing a planting strategy). The model must infer and a goal-state representation:
@@
 The generation should then follow:
-F.3 Evaluation Criteria
+G.3 Evaluation Criteria
 Predictions are evaluated against:
@@
 Goal Relevance
-F.4 Data Curation Details
+G.4 Data Curation Details
 Each task sample consists of the dialogue context, referenced images, and metadata such as source dialogue ID and topic.
@@
 Our modular pipeline supports deterministic regeneration of the dataset via fixed seeds and indexing, enabling reproducible experimentation and future extensibility to other domains.
-Release Protocol for MIRAGE-MMMT: In designing our dataset release, we follow established best practices from recent benchmarks such as MMLU [35], and BIG-Bench [32, 33], which emphasize the importance of separating training data and test targets to prevent leakage and ensure reliable model evaluation.
+Release Protocol for MIRAGE-MMMT: In designing our dataset release, we follow established best practices from recent benchmarks such as MMLU [hendrycks2020measuring], and BIG-Bench [ghazal2013bigbench, ghazal2017bigbench], which emphasize the importance of separating training data and test targets to prevent leakage and ensure reliable model evaluation.
 We adopt a protocol that maximizes transparency, reproducibility, and community usability, while preserving the integrity of the held-out test set.
@@
 Evaluation of model predictions on the test set can be conducted either via our LLM-based judge or via human assessment.
-F.5 Additional MIRAGE-MMMT Results
+G.5 Additional MIRAGE-MMMT Results
 | Input Variant | Decision Acc. | F1 (Macro) | Level | ||
@@
 | Dialogue + GoalState | 89.27% | 19.48% | 0.89 | 0.19 | Oracle |
-Appendix G Prompts
-G.1 Evaluation Prompts for MIRAGE-MMST
+Appendix H Prompts
+H.1 Evaluation Prompts for MIRAGE-MMST
 Figure 13 presents the inference prompt used for the MIRAGE-MMST Identification Task.
@@
 The field model_response refers to the generated answer being evaluated.
-G.2 Evaluation Prompts for MIRAGE-MMMT
-Appendix H Case Study
-[…]
-H.1 Category-Wise Cases
-H.1.1 Plant Identification (MMST Standard)
-H.1.2 Insect and Pest Identification (MMST Standard)
-H.1.3 Plant Disease Identification (MMST Standard)
-H.1.4 Plant Disease Management (MMST Standard)
-H.1.5 Insect and Pest Management (MMST Standard)
-H.1.6 Plant Care and Gardening Guidance (MMST Standard)
-H.1.7 Weeds/Invasive Plants Management (MMST Standard)
-H.1.8 Plant Disease Management (MMST Contextual)
-H.1.9 Insect and Pest Management (MMST Contextual)
-H.1.10 Plant Care and Gardening Guidance (MMST Contextual)
-H.1.11 Weeds/Invasive Plants Management (MMST Contextual)
-H.1.12 Clarify (MMMT)
-H.1.13 Respond (MMMT)
-H.2 Examples of Reasoning LLM as a Judge
-H.2.1 MMST Identification Task
-H.2.2 MMST Management Task
-H.2.3 MMMT Task
+H.2 Evaluation Prompts for MIRAGE-MMMT
+Appendix I Case Study
+I.1 Category-Wise Cases
+I.1.1 Plant Identification (MMST Standard)
+I.1.2 Insect and Pest Identification (MMST Standard)
+I.1.3 Plant Disease Identification (MMST Standard)
+I.1.4 Plant Disease Management (MMST Standard)
+I.1.5 Insect and Pest Management (MMST Standard)
+I.1.6 Plant Care and Gardening Guidance (MMST Standard)
+I.1.7 Weeds/Invasive Plants Management (MMST Standard)
+I.1.8 Plant Disease Management (MMST Contextual)
+I.1.9 Insect and Pest Management (MMST Contextual)
+I.1.10 Plant Care and Gardening Guidance (MMST Contextual)
+I.1.11 Weeds/Invasive Plants Management (MMST Contextual)
+I.1.12 Clarify (MMMT)
+I.1.13 Respond (MMMT)
+I.2 Examples of Reasoning LLM as a Judge
+I.2.1 MMST Identification Task
+I.2.2 MMST Management Task
+I.2.3 MMMT Task